### v1

# Phase transitions for unique codings of fat Sierpinski gasket with multiple digits

Abstract.
Given an integer $M\geq 1$ and $\beta\in(1,M+1)$, let $S_{\beta,M}$ be the fat Sierpinski gasket in $\mathbb{R}^{2}$ generated by the iterated function system $\left\{f_{d}(x)=\frac{x+d}{\beta}:d\in\Omega_{M}\right\}$, where $\Omega_{M}=\{(i,j)\in\mathbb{Z}_{\geq 0}^{2}:i+j\leq M\}$. Then each $x\in S_{\beta,M}$ may represent as a series $x=\sum_{i=1}^{\infty}\frac{d_{i}}{\beta^{i}}=:\Pi_{\beta}((d_{i}))$, and the infinite sequence $(d_{i})\in\Omega_{M}^{\mathbb{N}}$ is called a coding of $x$. Since $\beta<M+1$, a point in $S_{\beta,M}$ may have multiple codings. Let $U_{\beta,M}$ be the set of $x\in S_{\beta,M}$ having a unique coding, that is
$U_{\beta,M}=\left\{x\in S_{\beta,M}:\#\Pi_{\beta}^{-1}(x)=1\right\}.$
When $M=1$, Kong and Li [2020, Nonlinearity] described two critical bases for the phase transitions of the intrinsic univoque set $\widetilde{U}_{\beta,1}$, which is a subset of $U_{\beta,1}$. In this paper we consider $M\geq 2$, and characterize the two critical bases $\beta_{G}(M)$ and $\beta_{c}(M)$ for the phase transitions of $U_{\beta,M}$: (i) if $\beta\in(1,\beta_{G}(M)]$, then $U_{\beta,M}$ is finite; (ii) if $\beta\in(\beta_{G}(M),\beta_{c}(M))$ then $U_{\beta,M}$ is countably infinite; (iii) if $\beta=\beta_{c}(M)$ then $U_{\beta,M}$ is uncountable and has zero Hausdorff dimension; (iv) if $\beta>\beta_{c}(M)$ then $U_{\beta,M}$ has positive Hausdorff dimension. Our results can also be applied to the intrinsic univoque set $\widetilde{U}_{\beta,M}$. Moreover, we show that the first critical base $\beta_{G}(M)$ is a perron number, while the second critical base $\beta_{c}(M)$ is a transcendental number.
Key words and phrases:
Fat Sierpinski gasket, unique coding, phase transition, critical bases, Thue-Morse sequence.
2010 Mathematics Subject Classification:
Primary: 37B10, 28A80; Secondary: 68R15, 11A63
1. Introduction
Given $\beta>1$ and $M\in\mathbb{N}$, let $S_{\beta,M}$ be the Sierpinski gasket in $\mathbb{R}^{2}$ generated by the iterated function system (IFS) $\left\{f_{\beta,d}(x)=\frac{x+d}{\beta}:d\in\Omega_{M}\right\}$, where
$\Omega_{M}:=\left\{\alpha_{ij}=(i,j):i,j\in\left\{0,1,\ldots,M\right\}\textrm{and }i+j\leq M\right\}.$
It is well-known that $S_{\beta,M}$ is the unique nonempty compact set satisfying the equation $S_{\beta,M}=\bigcup_{d\in\Omega}f_{\beta,d}(S_{\beta,M})$ (cf. [18]). To study $S_{\beta,M}$ it is useful to introduce the coding map
(1.1) $\Pi_{\beta}:\Omega_{M}^{\mathbb{N}}\to S_{\beta,M};\quad(d_{i})\mapsto\sum_{i=1}^{\infty}\frac{d_{i}}{\beta^{i}},$
which is always continuous and surjective. If $\beta>M+1$, the IFS $\mathscr{F}_{\beta,M}=\left\{f_{\beta,d}:d\in\Omega_{M}\right\}$ satisfies the strong separation condition, and then $\Pi_{\beta}$ is bijective (cf. [15]). If $\beta=M+1$, the IFS $\mathscr{F}_{\beta,M}$ fails the strong separation condition but still admits the open set condition, and thus $\Pi_{\beta}$ is nearly bijective up to a countable set. However, for $\beta<M+1$ the IFS $\mathscr{F}_{\beta,M}$ fails the open set condition and it always has overlaps. In particular, if $\beta\in(1,\frac{M+2}{2}]$ the attractor $S_{\beta,M}$ is an isosceles triangle with vertices $(0,0),(\frac{M}{\beta-1},0)$ and $(0,\frac{M}{\beta-1})$ (see Proposition 2.5). In this case, for Lebesgue almost every $x\in S_{\beta,M}$ the preimage $\Pi_{\beta}^{-1}(x)$ is uncountable (cf. [28]). When $\beta\in(\frac{M+2}{2},M+1)$ the structure of $S_{\beta,M}$ is getting more complicated. Broomhead et al [5] showed that for $M=1$ and $\beta\leq\beta_{*}$ the set $S_{\beta,M}$ has non-empty interior, where $\beta_{*}\approx 1.54369$ is a root of $x^{3}-2x^{2}+2x-2=0$.
When $\beta\in(1,M+1)$ the fat Sierpinski gasket has attracted much attention in the past twenty years. Simon and Solomyak [29] showed that for $M=1$ there exists a dense set of $\beta\in(1,2)$ such that $\dim_{H}S_{\beta,1}<\log 3/\log\beta$, where $\log 3/\log\beta$ is the self-similar dimension of $S_{\beta,1}$. On the other hand, Jordan and Pollicott [19] proved that for $M=1$ the fat Sierpinski gasket $S_{\beta,1}$ has positive Lebesgue measure for Lebesgue almost every $\beta\in(1,1.70853]$. Recently, Hochman [17] showed that $\dim_{H}S_{\beta,M}=\min\left\{2,\dim_{S}S_{\beta,M}\right\}$ for $\beta\in(1,M+1)$ up to a set of zero packing dimension, where $\dim_{S}S_{\beta,M}=\frac{\log(M+1)+\log(M+2)-\log 2}{\log\beta}$ is the similarity dimension of $S_{\beta,M}$.
$x$$y$$\alpha_{00}$$\alpha_{10}$$\alpha_{20}$$\alpha_{30}$$\alpha_{40}$$\alpha_{01}$$\alpha_{02}$$\alpha_{03}$$\alpha_{04}$$\alpha_{11}$$\alpha_{12}$$\alpha_{13}$$\alpha_{21}$$\alpha_{31}$$\alpha_{22}$$\frac{1}{\beta}$$\frac{2}{\beta}$$\frac{3}{\beta}$$\frac{4}{\beta}$$\frac{4}{\beta-1}$$\frac{1}{\beta}$$\frac{2}{\beta}$$\frac{3}{\beta}$$\frac{4}{\beta}$$\frac{4}{\beta-1}$
Figure 1. The graph for the first generation of $S_{\beta,M}$ with $M=4$ and $\beta=4$. The convex hull $\Delta_{\beta,M}$ is the triangle with vertices $(0,0),(4/(\beta-1),0)$ and $(0,4/(\beta-1))$. Each light grey triangle corresponds to a $f_{\beta,\alpha_{ij}}(\Delta_{\beta,M})$ for some $\alpha_{ij}\in{\Omega}_{M}$, and the overlap region $O_{\beta,M}$ is the union of eighteen small dark grey triangles.
Observe that for $\beta\in(1,M+1)$ the overlap region
$O_{\beta,M}=\bigcup_{c,d\in{\Omega}_{M},c\neq d}f_{\beta,c}(\Delta_{\beta,M})\cap f_{\beta,d}(\Delta_{\beta,M})$
is nonempty, where $\Delta_{\beta,M}$ is the convex hull of $S_{\beta,M}$ (see Figure 1). Motivated by the work of Kong and Li [23], we considered the intrinsic univoque set
(1.2) $\widetilde{U}_{\beta,M}:=\left\{\Pi_{\beta}((d_{i}))\in S_{\beta,M}:\Pi_{\beta}(d_{n+1}d_{n+2}\ldots)\notin O_{\beta,M}\leavevmode\nobreak\ \forall n\geq 0\right\},$
where $\Pi_{\beta}$ is defined in (1.1). Note that $\widetilde{U}_{\beta,M}$ is the survivor set of some open dynamical systems (cf. [27]). Indeed, let $T_{\beta,M}$ be the expanding map from the isosceles triangle $\Delta_{\beta,M}$ to itself such that $T_{\beta,M}(x)=f_{\beta,d}^{-1}(x)$ if $x\in f_{\beta,d}(\Delta_{\beta,M})$ for any $d\in{\Omega}_{M}$. Then $S_{\beta,M}$ is invariant under $T_{\beta,M}$, and the intrinsic univoque set $\widetilde{U}_{\beta,M}$ is the survivor set consisting of all $x\in\Delta_{\beta,M}$ whose orbit under $T_{\beta,M}$ never enters the hole $O_{\beta,M}$, i.e., $\widetilde{U}_{\beta,M}=\left\{x\in\Delta_{\beta,M}:T_{\beta,M}^{n}(x)\notin O_{\beta,M}\leavevmode\nobreak\ \forall n\geq 0\right\}$. For more information on open dynamical systems we refer to the papers [30, 12, 13, 11, 16, 7, 6, 20, 1] and the references therein.
Since $\beta\in(1,M+1)$, $S_{\beta,M}$ is a self-similar set with overlaps. Inspired by the work of unique $\beta$-expansions (cf. [14, 9, 21]), we study the univoque set $U_{\beta,M}$ in $S_{\beta,M}$, i.e.,
(1.3) $U_{\beta,M}=\left\{x\in S_{\beta,M}:\#\Pi_{\beta}^{-1}(x)=1\right\},$
where #$\#A$ denotes the cardinality of a set $A$. Note that each point in $\widetilde{U}_{\beta,M}$ has a unique coding under $\Pi_{\beta}$. Then $\widetilde{U}_{\beta,M}\subset U_{\beta,M}$ for any $\beta\in(1,M+1)$.
When $\beta\in(1,M+1)$ is getting larger, the overlap region $O_{\beta,M}$ is getting relatively smaller, and it turns out that $\widetilde{U}_{\beta,M}$ is getting larger. In this paper we will describe the phase transition phenomenon for $\widetilde{U}_{\beta,M}$ and $U_{\beta,M}$. More precisely, we will describe two critical bases $\beta_{G}(M)$ and $\beta_{c}(M)$ with $\beta_{G}(M)<\beta_{c}(M)$. The first critical base $\beta_{G}(M)$, called the generalized golden ratio, is the unique base such that the cardinality of $\widetilde{U}_{\beta,M}$ increases from finite to infinite when $\beta$ passes $\beta_{G}(M)$. The second critical base $\beta_{c}(M)$, called the generalized Komornik-Loreti constant, is the unique base in which the Hausdorff dimension of $\widetilde{U}_{\beta,M}$ changes from zero to positive when $\beta$ crosses $\beta_{c}(M)$. When $M\geq 2$, we will show that the two critical bases $\beta_{G}(M)$ and $\beta_{c}(M)$ can also be applied to $U_{\beta,M}$.
Denote by :=$\mathbb{N}:=\left\{1,2,3,\ldots\right\}$ and $\mathbb{N}_{0}:=\mathbb{N}\cup\left\{0\right\}$. Let $(\tau_{i})_{i=0}^{\infty}=01101001\ldots\in\left\{0,1\right\}^{\mathbb{N}}$ be the classical Thue-Morse sequence defined by (cf. [2])
$\tau_{0}=0,\quad\tau_{2i+1}=1-\tau_{i}\quad\textrm{and}\quad\tau_{2i}=\tau_{i}\quad\forall i\geq 0.$
Now we are ready to define the two critical bases $\beta_{G}(M)$ and $\beta_{c}(M)$.
Definition 1.1.
The generalized golden ratio $\beta_{G}(M)$ for $M\geq 1$ is defined as follows. If $M=3N+1$ for some $N\in\mathbb{N}_{0}:=\left\{0,1,2,\ldots\right\}$, then $\beta_{G}(3N+1)$ is the unique root in $(N+1,N+2)$ of the equation
$x^{3}-(N+1)x^{2}-Nx-(N+1)=0.$
Similarly, if $M=3N+2$ with $N\in\mathbb{N}_{0}$, then $\beta_{G}(3N+2)$ is the unique root in $(N+1,N+2)$ of the equation
$x^{3}-(N+1)(x^{2}+x+1)=0.$
Finally, if $M=3N+3$ with $N\in\mathbb{N}_{0}$, then $\beta_{G}(3N+3)=N+2$.
Definition 1.2.
The generalized Komornik-Loreti constant $\beta_{c}(M)$ for $M\geq 1$ is defined as follows. If $M=3N+1$ for some $N\in\mathbb{N}_{0}$, then $\beta_{c}(3N+1)$ is the unique root in $(N+1,N+2)$ of the equation
$\sum_{i=0}^{\infty}\left(\frac{N+\tau_{2i+1}}{x^{3i+1}}+\frac{N}{x^{3i+2}}+\frac{N+\tau_{2i+2}}{x^{3i+3}}\right)=1,$
where $(\tau_{i})_{i=0}^{\infty}=01101001\ldots$ is the classical Thue-Morese sequence. Similarly, if $M=3N+2$ with $N\in\mathbb{N}_{0}$, then $\beta_{c}(3N+2)$ is the unique root in $(N+1,N+2)$ of the equation
$\sum_{i=0}^{\infty}\left(\frac{N+\tau_{2i+1}}{x^{3i+1}}+\frac{N+1}{x^{3i+2}}+\frac{N+\tau_{2i+2}}{x^{3i+3}}\right)=1.$
Finally, if $M=3N+3$ with $N\in\mathbb{N}_{0}$, then $\beta_{c}(3N+3)$ is the unique root in $(N+2,N+3)$ of the equation
$\sum_{i=0}^{\infty}\left(\frac{N+2\tau_{2i+1}}{x^{3i+1}}+\frac{N+\tau_{2i}}{x^{3i+2}}+\frac{N+1+\tau_{2i+2}}{x^{3i+3}}\right)=1.$
By Definitions 1.1 and 1.2 it follows that
$[\frac{M}{3}]+1\leq\beta_{G}(M)<\beta_{c}(M)<[\frac{M}{3}]+2,$
where $[r]$ denotes the integer part of a real number $r$. Then $\lim_{M\to\infty}\frac{\beta_{G}(M)}{M}=\lim_{M\to\infty}\frac{\beta_{c}(M)}{M}=\frac{1}{3}$. In the following table we numerically calculate the two critical bases $\beta_{G}(M)$ and $\beta_{c}(M)$ for $M=1,2,\ldots,10$; and they are also illustrated in Figure 2.
$M$ 1 2 3 4 5 6 7 8 9 10
$\beta_{G}(M)\approx$ 1.46557 1.83929 2 2.65897 2.91964 3 3.74735 3.95137 4 4.79885
$\beta_{c}(M)\approx$ 1.55356 1.91988 2.40462 2.72236 2.97737 3.55447 3.78826 3.98799 4.64302 4.82717
Figure 2. The graph of the two critical bases $\beta_{G}(M)$ and $\beta_{c}(M)$ with $M=1,2,\ldots,10$.
We will show that each $\beta_{G}(M)$ is a Perron number, and each $\beta_{c}(M)$ is a transcendental number. When $M=1$ the following phase transition phenomenon for $\widetilde{U}_{\beta,1}$ was established by Sidorov [28, Theorem 4.1] and Kong and Li [23, Theorem 2].
Theorem 1.3 ([28, 23]).
Let $\beta_{G}(1)\approx 1.46557$ and $\beta_{c}(1)\approx 1.55356$ be defined as in Definitions 1.1 and 1.2, respectively. Then $\beta_{G}(1)$ is a Perron number, and $\beta_{c}(1)$ is a transcendental number.
  1. (i)
    If $\beta\in(1,\beta_{G}(1)]$, then #$\#\widetilde{U}_{\beta,1}=\#U_{\beta,1}=3$;
  2. (ii)
    If $\beta\in(\beta_{G}(1),\beta_{c}(1))$, then $\widetilde{U}_{\beta,1}$ is countably infinite;
  3. (iii)
    If $\beta=\beta_{c}(1)$, then $\widetilde{U}_{\beta,1}$ is uncountable and has zero Hausdorff dimension;
  4. (iv)
    If $\beta\in(\beta_{c}(1),2)$, then $\widetilde{U}_{\beta,1}$ has positive Hausdorff dimension.
Since $\widetilde{U}_{\beta,1}\subset U_{\beta,1}$, by Theorem 1.3 it follows that $U_{\beta,1}$ is infinite for $\beta>\beta_{G}(1)$. So, Theorem 1.3 implies that $\beta_{G}(1)$ is also a critical base for $U_{\beta,1}$, i.e., $U_{\beta,1}$ is finite if and only if $\beta\leq\beta_{G}(1)$. However, we don’t know whether $\beta_{c}(1)$ is still the critical base for $U_{\beta,1}$ in the sense that $U_{\beta,1}$ has positive Hausdorff dimension if and only if $\beta>\beta_{c}(1)$.
In this paper we extend Theorem 1.3 to $M\geq 2$, and show that (i) $\widetilde{U}_{\beta,M}$ and $U_{\beta,M}$ are finite if and only if $\beta\leq\beta_{G}(M)$; (ii) $\widetilde{U}_{\beta,M}$ and $U_{\beta,M}$ have positive Hausdorff dimension if and only if $\beta>\beta_{c}(M)$. The new challenge is that for $M=3N+1$ with $N\geq 1$ the digit $N-1$ might occur in $d_{n}^{1}$ or $d_{n}^{2}$ for a sequence $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{c},M}$, while for $M=1$ this can never happen. Fortunately, we are able to show that this digit $N-1$ can occur at most 3 times in any sequence of $\widetilde{{\mathbf{U}}}_{\beta_{c},M}$ (see Lemma 4.5). Furthermore, completely different from the case for $M=1$, we need more effort to find the critical base $\beta_{c}(M)$ for $M=3N+3$. In other words, we need to construct a completely new type of sequences to represent $\beta_{c}(M)$, see more in Section 6.
Now we state the main result of the paper.
Theorem 1.4.
Given $M\geq 2$, let $\beta_{G}(M)$ and $\beta_{c}(M)$ be defined as in Definitions 1.1 and 1.2 respectively. Then $\beta_{G}(M)$ is a Perron number, and $\beta_{c}(M)$ is a transcendental number.
  1. (i)
    If $\beta\in(1,\beta_{G}(M)]$, then #$\#U_{\beta,M}=\#\widetilde{U}_{\beta,M}=3$;
  2. (ii)
    If $\beta\in(\beta_{G}(M),\beta_{c}(M))$, then $U_{\beta,M}=\widetilde{U}_{\beta,M}$ is countably infinite;
  3. (iii)
    If $\beta=\beta_{c}(M)$, then $U_{\beta,M}=\widetilde{U}_{\beta,M}$ is uncountable and has zero Hausdorff dimension;
  4. (iv)
    If $\beta>\beta_{c}(M)$, then both $U_{\beta,M}$ and $\widetilde{U}_{\beta,M}$ have positive Hausdorff dimensions.
When $\beta>\beta_{c}(M)$ and $\beta$ is a generalized multinacci number, we are able to give the explicit formula for the Hausdorff dimension of $U_{\beta,M}$ (see Example 4.13 below).
The rest of the paper is organized as follows. In the next section we give the characterization of $\widetilde{U}_{\beta,M}$. In Section 3 we show that the generalized golden ratio $\beta_{G}(M)$ defined in Definition 1.1 is the first critical base for $\widetilde{U}_{\beta,M}$ that $\widetilde{U}_{\beta,M}$ is finite if and only if $\beta\leq\beta_{G}(M)$. Furthermore, we show that the generalized Komornik-Loreti constant $\beta_{c}(M)$ in Definition 1.2 is the second critical base for $\widetilde{U}_{\beta,M}$. This will be proved in Sections 46 for $M=3N+1,M=3N+2$ and $M=3N+3$, respectively. We point out that the two critical bases $\beta_{G}(M)$ and $\beta_{c}(M)$ are also applied to $U_{\beta,M}$ by Proposition 2.5.
2. Characterization of $\widetilde{U}_{\beta,M}$
Given $M\in\mathbb{N}$, we recall ${\Omega}_{M}=\left\{\alpha_{ij}=(i,j)\in\mathbb{N}^{2}_{0}:i+j\leq M\right\}.$ For $d=(d^{1},d^{2})\in\Omega_{M}$ let $d^{\oplus}:=d^{1}+d^{2}$ and $\overline{d^{\oplus}}:=M-d^{\oplus}$. Then
(2.1) $d^{1}+d^{2}+\overline{d^{\oplus}}=M\quad\textrm{for any }d\in\Omega_{M}.$
First we introduce some terminology from symbolic dynamics (cf. [24]). For $n\in\mathbb{N}$ let ${\Omega}_{M}^{n}$ be the set of all length $n$ words with each digit from ${\Omega}_{M}$. Denote by ${\Omega}_{M}^{*}:=\bigcup_{n=1}^{\infty}{\Omega}_{M}^{n}$ the set of all finite words over the alphabet ${\Omega}_{M}$. Moreover, let ${\Omega}_{M}^{\mathbb{N}}$ be the set of all infinite sequences with each digit from ${\Omega}_{M}$. For a word $d_{1}d_{2}\ldots d_{n}\in{\Omega}_{M}^{*}$ or a sequence $(d_{i})\in{\Omega}_{M}^{\mathbb{N}}$ it is useful to represent it in a matrix form:
$d_{1}d_{2}\ldots d_{n}\sim\left(\begin{array}[]{c}d_{1}^{1}d_{2}^{1}\ldots d_{n}^{1}\\ d_{1}^{2}d_{2}^{2}\ldots d_{n}^{2}\\ \overline{d_{1}^{\oplus}}\,\overline{d_{2}^{\oplus}}\ldots\overline{d_{n}^{\oplus}}\end{array}\right),\quad d_{1}d_{2}\ldots\sim\left(\begin{array}[]{c}d_{1}^{1}d_{2}^{1}\ldots\\ d_{1}^{2}d_{2}^{2}\ldots\\ \overline{d_{1}^{\oplus}}\,\overline{d_{2}^{\oplus}}\ldots\end{array}\right).$
This representation plays an important role in our proof of Theorem 1.4; and it will be clear after we proving Proposition 2.3. Note by (2.1) that each row in the above presentation is a length $n$ word or an infinite sequence over the alphabet $\left\{0,1,\ldots,M\right\}$.
Let $\theta:{\Omega}_{M}\to{\Omega}_{M}$ be a cyclic permutation map defined as follows. For $d=(d^{1},d^{2})\in\Omega_{M}$ set
(2.2) $\theta(d)=\theta((d^{1},d^{2}))=(M-(d^{1}+d^{2}),d^{1})=(\overline{d^{\oplus}},d^{1}).$
Then $\theta({\Omega}_{M})={\Omega}_{M}$. Furthermore, $\theta^{2}(d)=(d^{2},\overline{d^{\oplus}})$ and $\theta^{3}(d)=d$. In other words, by using the matrix form we have
$\left(\begin{array}[]{c}d^{1}\\ d^{2}\\ \overline{d^{\oplus}}\end{array}\right)\stackrel{{\scriptstyle\theta}}{{\longrightarrow}}\left(\begin{array}[]{c}\overline{d^{\oplus}}\\ d^{1}\\ d^{2}\end{array}\right)\stackrel{{\scriptstyle\theta^{2}}}{{\longrightarrow}}\left(\begin{array}[]{c}d^{2}\\ \overline{d^{\oplus}}\\ d^{1}\end{array}\right)\stackrel{{\scriptstyle\theta^{3}}}{{\longrightarrow}}\left(\begin{array}[]{c}d^{1}\\ d^{2}\\ \overline{d^{\oplus}}\end{array}\right).$
For a block $d_{1}\ldots d_{n}\in{\Omega}_{M}^{*}$ we define $\theta(d_{1}\ldots d_{n})=\theta(d_{1})\ldots\theta(d_{n})\in{\Omega}_{M}^{*}$; and for a sequence $(d_{i})\in{\Omega}_{M}^{\mathbb{N}}$ we set $\theta((d_{i}))=\theta(d_{1})\theta(d_{2})\ldots\in{\Omega}_{M}^{\mathbb{N}}$.
Given $\beta\in(1,M+1)$, for an infinite sequence $(d_{i})\in\Omega_{M}^{\mathbb{N}}$ we write
$\Pi_{\beta}(d_{1}d_{2}\ldots)=\sum_{i=1}^{\infty}\frac{d_{i}}{\beta^{i}}=\left(\sum_{i=1}^{\infty}\frac{d_{i}^{1}}{\beta^{i}},\sum_{i=1}^{\infty}\frac{d_{i}^{2}}{\beta^{i}}\right)=:(\pi_{\beta}((d_{i}^{1})),\pi_{\beta}((d_{i}^{2}))).$
Note that the two induced sequences $(d_{i}^{1})$ and $(d_{i}^{2})$ both have digits in $\left\{0,1,\ldots,M\right\}$. Then we also need to introduce the one-dimensional symbolic dynamics ($(\left\{0,1,\ldots,M\right\}^{\mathbb{N}},\sigma)$, where $\left\{0,1,\ldots,M\right\}^{\mathbb{N}}$ is the set of all infinite sequences with each digit from $\left\{0,1,\ldots,M\right\}$, and $\sigma$ is the left shift map. For $n\in\mathbb{N}$ let $\left\{0,1,\ldots,M\right\}^{n}$ be the set of all length $n$ words with each digit from $\left\{0,1,\ldots,M\right\}$. Set $\left\{0,1,\ldots,M\right\}^{*}:=\bigcup_{n=1}^{\infty}\left\{0,1,\ldots,M\right\}^{n}$. For a word $w=w_{1}\ldots w_{n}\in\left\{0,1,\ldots,M\right\}^{*}$, if $w_{n}<M$ then we set $w^{+}:=w_{1}\ldots w_{n-1}(w_{n}+1)$; if $w_{n}>0$ then we set $w^{-}:=w_{1}\ldots w_{n-1}(w_{n}-1)$. We also write $w^{\infty}$ for the periodic sequence with periodic block $w$. For a sequence $(c_{i})\in\left\{0,1,\ldots,M\right\}^{\mathbb{N}}$ we denote its reflection by $\overline{(c_{i})}=(M-c_{1})(M-c_{2})\ldots$. Furthermore, we use lexicographical order between sequences in $\left\{0,1,\ldots,M\right\}^{\mathbb{N}}$. For two infinite sequences $(c_{i}),(d_{i})\in\left\{0,1,\ldots,M\right\}^{\mathbb{N}}$ we write $(c_{i})\prec(d_{i})$ if there exists $N\in\mathbb{N}$ such that $c_{1}\ldots c_{N-1}=d_{1}\ldots d_{N-1}$ and $c_{N}<d_{N}$. We also write $(c_{i})\preccurlyeq(d_{i})$ if $(c_{i})\prec(d_{i})$ or $(c_{i})=(d_{i})$. Similarly, we write $(c_{i})\succ(d_{i})$ if $(d_{i})\prec(c_{i})$; and write $(c_{i})\succcurlyeq(d_{i})$ if $(d_{i})\preccurlyeq(c_{i})$.
For $\beta\in(1,M+1)$ let $\delta(\beta)=\delta_{1}(\beta)\delta_{2}(\beta)\ldots\in\left\{0,1,\ldots,M\right\}^{\mathbb{N}}$ be the quasi-greedy $\beta$-expansion of $1$, i.e., the lexicographically largest sequence $(\delta_{i})$ in $\left\{0,1,\ldots,M\right\}^{\mathbb{N}}$ not ending with $0^{\infty}$ and satisfying $\pi_{\beta}((\delta_{i}))=1$ (cf. [8]). The following characterization of $\delta(\beta)$ was proven in [4].
Lemma 2.1.
The map $\beta\mapsto\delta(\beta)$ is strictly increasing and bijective from $(1,M+1]$ to the set of sequences $(\delta_{i})\in\{0,1,\ldots,M\}^{\mathbb{N}}$ satisfying
$0^{\infty}\prec\sigma^{n}((\delta_{i}))\preccurlyeq(\delta_{i})\quad\forall n\geq 0$
Furthermore, the inverse map $\delta^{-1}$ is continuous with respect to the order topology.
In general, for $\beta\in(1,M+1)$ and $x\in(0,1]$ let $a(x,\beta)$ be the quasi-greedy $\beta$-expansion of $x$, i.e., the lexicographically largest sequence $(a_{i})\in\left\{0,1,\ldots,M\right\}^{\mathbb{N}}$ not ending with $0^{\infty}$ and satisfying $\pi_{\beta}((a_{i}))=x$. The following result for quasi-greedy expansions was established by Parry [26] and de Vries, Komornik and Loreti [10, Lemma 2.2 and Proposition 2.4].
Lemma 2.2.
Let $\beta\in(1,M+1]$. The map $x\mapsto a(x,\beta)$ is strictly increasing and bijective from $(0,1]$ to the set
$\left\{(a_{i})\in\left\{0,1,\ldots,M\right\}^{\mathbb{N}}:0^{\infty}\prec\sigma^{n}((a_{i}))\preccurlyeq\delta(\beta)\leavevmode\nobreak\ \forall n\geq 0\right\}.$
Furthermore, the map $x\mapsto a(x,\beta)$ is left-continuous with respect to the order topology.
For $\beta\in(1,M+1)$ we recall from (1.2) and (1.3) the definitions of intrinsic univoque set $\widetilde{U}_{\beta,M}$ and univoque set ${U}_{\beta,M}$, respectively. According to the coding map $\Pi_{\beta}$ in (1.1), we define the symbolic intrinsic univoque set $\widetilde{{\mathbf{U}}}_{\beta,M}\subset{\Omega}_{M}^{\mathbb{N}}$ and symbolic univoque set ${\mathbf{U}}_{\beta,M}\subset{\Omega}_{M}^{\mathbb{N}}$ by
$\widetilde{{\mathbf{U}}}_{\beta,M}:=\Pi_{\beta}^{-1}(\widetilde{U}_{\beta,M})\quad\textrm{and}\quad{\mathbf{U}}_{\beta,M}:=\Pi_{\beta}^{-1}(U_{\beta,M}).$
Note that $\widetilde{U}_{\beta,M}\subset U_{\beta,M}$ for any $\beta\in(1,M+1)$, and each $x\in U_{\beta,M}$ has a unique coding under $\Pi_{\beta}$. Then $\Pi_{\beta}$ is bijective from $\widetilde{{\mathbf{U}}}_{\beta,M}$ to $\widetilde{U}_{\beta,M}$, and from ${\mathbf{U}}_{\beta,M}$ to $U_{\beta,M}$ as well.
The following characterization of $\widetilde{{\mathbf{U}}}_{\beta,M}$ plays an important role in the proof of Theorem 1.4. When $M=1$, it was proven in [23, Proposition 2.2].
Proposition 2.3.
Let $M\in\mathbb{N}$ and $\beta\in(1,M+1)$. Then $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta,M}$ if and only if $(d_{i})\in\Omega_{M}^{\mathbb{N}}$ satisfies
(2.3) $\left\{\begin{array}[]{ccc}d_{n+1}^{1}d_{n+2}^{1}\ldots\prec\delta(\beta)&\textrm{whenever}&d_{n}^{1}<M,\\ d_{n+1}^{2}d_{n+2}^{2}\ldots\prec\delta(\beta)&\textrm{whenever}&d_{n}^{2}<M,\\ \overline{d_{n+1}^{\oplus}}\overline{d_{n+2}^{\oplus}}\ldots\prec\delta(\beta)&\textrm{whenever}&\overline{d_{n}^{\oplus}}<M.\end{array}\right.$
Proof.
First we prove the necessity. Take $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta,M}$. We will prove (2.3) by considering the following three cases: (I) $d_{n}^{1}<M$; (II) $d_{n}^{2}<M$; (III) $\overline{d_{n}^{\oplus}}<M$ for some $n\in\mathbb{N}$.
$\frac{s}{\beta}$$\frac{s+t-1}{\beta}+\frac{M}{\beta(\beta-1)}$$\frac{s+1}{\beta}$$\frac{s+t}{\beta}+\frac{M}{\beta(\beta-1)}$$\frac{t}{\beta}$$\frac{t+1}{\beta}$$\alpha_{st}$
Figure 3. The typical pattern of $f_{\beta,d}(\Delta_{\beta,M})\setminus O_{\beta,M}$ with $d=(d^{1},d^{2})\in\Omega_{M}$ satisfying $d^{1}<M,d^{2}<M$ and $\overline{d^{\oplus}}=M-(d^{1}+d^{2})<M$.
Case (I). Suppose $d_{n}^{1}=s<M$ for some $n$. Then $d_{n}=\alpha_{sj}$ for some $j\in\left\{0,1,\ldots,M\right\}$. Since $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta,M}=\Pi_{\beta}^{-1}(\widetilde{U}_{\beta,M})$, by the definition of $\widetilde{U}_{\beta,M}$ in (1.3) it follows that
$\Pi_{\beta}(d_{n}d_{n+1}\ldots)\in\bigcup_{j=0}^{M}f_{\beta,\alpha_{sj}}(\Delta_{\beta,M})\setminus{O}_{\beta,M}.$
In view of Figure 1, there are four possible patterns for $f_{\beta,d}(\Delta_{\beta,M})\setminus O_{\beta,M}$ with $d\in\Omega_{M}$: see Figure 3 for the typical pattern, and see Figure LABEL:fig:1' for the patterns in the three corners. Since $d_{n}^{1}=s<M$, this implies $\pi_{\beta}(d_{n}^{1}d_{n+1}^{1}\ldots)<\frac{s+1}{\beta}$. Note that $d_{n}^{1}=s$. Then
(2.4) $\pi_{\beta}(d_{n+1}^{1}d_{n+2}^{1}\ldots)<1=\pi_{\beta}(\delta(\beta)).$
So, it suffices to prove $d_{n+1}^{1}d_{n+2}^{1}\ldots\prec\delta(\beta)$.
Suppose on the contrary that $d_{n+1}^{1}d_{n+2}^{1}\ldots\succcurlyeq\delta(\beta)$. If the equality holds, then we have $\pi_{\beta}(d_{n+1}^{1}d_{n+2}^{1}\ldots)=1$, leading to a contradiction with (2.4). If strict inequality holds, then there exists $k\in\mathbb{N}$ such that $d_{n+1}^{1}\ldots d_{n+k-1}^{1}=\delta_{1}(\beta)\ldots\delta_{k-1}(\beta)$ and $d_{n+k}^{1}>\delta_{k}(\beta)$. By Lemmas 2.1 and 2.2 it follows
$1=\pi_{\beta}(\delta(\beta))$ $=\sum_{i=1}^{k}\frac{\delta_{i}(\beta)}{\beta^{i}}+\frac{1}{\beta^{k}}\sum_{i=1}^{\infty}\frac{\delta_{k+i}(\beta)}{\beta^{i}}$
$\leq\sum_{i=1}^{k}\frac{\delta_{i}(\beta)}{\beta^{i}}+\frac{1}{\beta^{k}}$
$\leq\sum_{i=1}^{k}\frac{d_{n+i}^{1}}{\beta^{i}}\leq\pi_{\beta}(d_{n+1}^{1}d_{n+2}^{1}\ldots),$
again leading to a contradiction with (2.4). This proves $d_{n+1}^{1}d_{n+2}^{1}\ldots\prec\delta(\beta)$ as required.
Case (II). Suppose $d_{n}^{2}=t<M$ for some $n$. Similar to Case (I) we can prove that $d_{n+1}^{2}d_{n+2}^{2}\ldots\prec\delta(\beta)$.
Case (III). Suppose $\overline{d_{n}^{\oplus}}=M-(s+t)<M$ for some $n$, where $d_{n}^{1}=s$ and $d_{n}^{2}=t$. Then $d_{n}^{\oplus}=d_{n}^{1}+d_{n}^{2}=M-\overline{d_{n}^{\oplus}}=s+t>0$. Since $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta,M}=\Pi_{\beta}^{-1}(\widetilde{U}_{\beta,M})$, by (1.3) we obtain that (see Figure 3)
$\pi_{\beta}(d_{n}^{\oplus}d_{n+1}^{\oplus}\ldots)=\pi_{\beta}(d_{n}^{1}d_{n+1}^{1}\ldots)+\pi_{\beta}(d_{n}^{2}d_{n+1}^{2}\ldots)>\frac{s+t-1}{\beta}+\frac{M}{\beta(\beta-1)}.$
Note that $d_{n}^{\oplus}=d_{n}^{1}+d_{n}^{2}=s+t$. Then $\pi_{\beta}(d_{n+1}^{\oplus}d_{n+2}^{\oplus}\ldots)>\frac{M}{\beta-1}-1,$ which implies
$\pi_{\beta}(\overline{d_{n+1}^{\oplus}d_{n+2}^{\oplus}\ldots})<1=\pi_{\beta}(\delta(\beta)).$
By the same argument as in Case (I) we conclude $\overline{d_{n+1}^{\oplus}d_{n+2}^{\oplus}}\prec\delta(\beta)$.
Next we prove the sufficiency. Let $(d_{i})\in\Omega_{M}^{\mathbb{N}}$ be an infinite sequence satisfying (2.3). Then it suffices to prove
(2.5) $\Pi_{\beta}(d_{n}d_{n+1}\ldots)\in f_{\beta,d_{n}}(\Delta_{\beta,M})\setminus O_{\beta,M}\quad\forall n\geq 1.$
Take $d_{n}=(s,t)\in{\Omega}_{M}$ for some $n\in\mathbb{N}$. In view of Figures 1 and 3, we split the proof of (2.5) into the following four cases.
Case A. $s<M,t<M$ and $s+t>0$. Then $d_{n}^{1}=s<M,d_{n}^{2}=t<M$ and $\overline{d_{n}^{\oplus}}=M-(s+t)<M$. First we claim that
(2.6) $\pi_{\beta}(d_{n}^{1}d_{n+1}^{1}\ldots)<\frac{s+1}{\beta}.$
Note that $d_{n}^{1}=s<M$. Starting with $k_{0}=n$ we define by (2.3) an infinite sequence of indices $k_{0}<k_{1}<k_{2}<\cdots$ satisfying for $j=1,2,\ldots$ the conditions
$d_{k_{j-1}+1}^{1}\ldots d_{k_{j}-1}^{1}=\delta_{1}(\beta)\ldots\delta_{k_{j}-k_{j-1}-1}(\beta)\quad\textrm{and}\quad d_{k_{j}}^{1}<\delta_{k_{j}-k_{j-1}}(\beta).$
Then by Lemma 2.1 it follows that
$\pi_{\beta}(d_{n}^{1}d_{n+1}^{1}\ldots)$ $=\frac{s}{\beta}+\beta^{n-1}\sum_{i=n+1}^{\infty}\frac{d_{i}^{1}}{\beta^{i}}=\frac{s}{\beta}+\beta^{n-1}\sum_{j=1}^{\infty}\sum_{i=1}^{k_{j}-k_{j-1}}\frac{d_{k_{j-1}+i}^{1}}{\beta^{k_{j-1}+i}}$
$\leq$ $\frac{s}{\beta}+\beta^{n-1}\sum_{j=1}^{\infty}\left(\sum_{i=1}^{k_{j}-k_{j-1}}\frac{\delta_{i}(\beta)}{\beta^{k_{j-1}+i}}-\frac{1}{\beta^{k_{j}}}\right)$
$<$ $\frac{s}{\beta}+\beta^{n-1}\sum_{j=1}^{\infty}\left(\frac{1}{\beta^{k_{j-1}}}-\frac{1}{\beta^{k_{j}}}\right)=\frac{s}{\beta}+\frac{\beta^{n-1}}{\beta^{k_{0}}}=\frac{s+1}{\beta}.$
This proves (2.6). Similarly, since $d_{n}^{2}=t<M$ and $\overline{d_{n}^{\oplus}}=M-(s+t)<M$, by (2.3) and the same argument as above we can prove
$\pi_{\beta}(d_{n}^{2}d_{n+1}^{2}\ldots)<\frac{t+1}{\beta}\quad\textrm{and}\quad\pi_{\beta}(\overline{d_{n}^{\oplus}d_{n+1}^{\oplus}\ldots})<\frac{s+t}{\beta};$
and the second inequality implies $\pi_{\beta}(d_{n}^{\oplus}d_{n+1}^{\oplus}\ldots)>\frac{s+t-1}{\beta}+\frac{M}{\beta(\beta-1)}$. Therefore, in view of Figure 3, this proves (2.5).
[…]
Figure 4. The three patterns of $f_{\beta,d}(\Delta_{\beta,M})\setminus O_{\beta,M}$ with $d=(0,0),d=(M,0)$ and $d=(0,M)$ respectively.
Case B. $s+t=0$. Then $d_{n}^{1}=d_{n}^{2}=0$ and $\overline{d_{n}^{\oplus}}=M$. By the same argument as in Case A we can prove that
$\pi_{\beta}(d_{n}^{1}d_{n+1}^{1}\ldots)<\frac{1}{\beta},\quad\pi_{\beta}(d_{n}^{2}d_{n+1}^{2}\ldots)<\frac{1}{\beta}.$
In view of Figure 4 (left graph), we can also prove (2.5).
Case C. $s=M$. Then $d_{n}^{1}=M$, $d_{n}^{2}=0$ and $\overline{d_{n}^{\oplus}}=0$. Similar to Case A, we have
$\pi_{\beta}(d_{n}^{2}d_{n+1}^{2}\ldots)<\frac{1}{\beta},\quad\pi_{\beta}(d_{n}^{\oplus}d_{n+1}^{\oplus}\ldots)>\frac{M-1}{\beta}+\frac{M}{\beta(\beta-1)}.$
In view of Figure 4 (middle graph), this yields (2.5).
Case D. $t=M$. Then $d_{n}^{1}=0$, $d_{n}^{2}=M$ and $\overline{d_{n}^{\oplus}}=0$. By the same argument as in Case A we obtain that
$\pi_{\beta}(d_{n}^{1}d_{n+1}^{1}\ldots)<\frac{1}{\beta},\quad\pi_{\beta}(d_{n}^{\oplus}d_{n+1}^{\oplus}\ldots)>\frac{M-1}{\beta}+\frac{M}{\beta(\beta-1)}.$
In view of Figure 4 (right graph), this also proves (2.5). ∎
Remark 2.4.
Note by (2.2) that $\theta$ is the cyclic permutation. Then by Proposition 2.3 it follows that
(2.7) $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta,M}\quad\Longleftrightarrow\quad\theta((d_{i}))\in\widetilde{{\mathbf{U}}}_{\beta,M}\quad\Longleftrightarrow\quad\theta^{2}((d_{i}))\in\widetilde{{\mathbf{U}}}_{\beta,M}.$
At the end of this section we show that for $\beta\leq(M+2)/2$ the univoque set ${\mathbf{U}}_{\beta,M}$ coincides with the intrinsic univoque set $\widetilde{{\mathbf{U}}}_{\beta,M}$.
Proposition 2.5.
Let $M\in\mathbb{N}$ and $\beta\in(1,M+1)$. Then $S_{\beta,M}=\Delta_{\beta,M}$ if and only if $\beta\in(1,\frac{M+2}{2}]$. Hence, for any $\beta\in(1,\frac{M+2}{2}]$ we have
$U_{\beta,M}=\widetilde{U}_{\beta,M}\quad\textrm{and}\quad{\mathbf{U}}_{\beta,M}=\widetilde{{\mathbf{U}}}_{\beta,M}.$
Proof.
In view of Figures 1 and 3, the fat Sierpinski gasket $S_{\beta,M}$ is equal to its convex hull $\Delta_{\beta,M}$ if and only if for any $d=(s,t)\in\mathbb{N}_{0}^{2}$ with $s+t\leq M$ we must have
$\frac{s+1}{\beta}+\frac{t+1}{\beta}\leq\frac{s+t}{\beta}+\frac{M}{\beta(\beta-1)},$
which is equivalent to $\beta\leq\frac{M+2}{2}$. This proves the first statement.
Suppose $\beta\in(1,\frac{M+2}{2}]$. Then $S_{\beta,M}=\Delta_{\beta,M}$, and in this case each point in the overlap region $O_{\beta,M}$ has at least two codings (cf. [23, Proposition 2.4]). So, by (1.2) and (1.3) it follows that $\widetilde{U}_{\beta,M}=U_{\beta,M}$, and also $\widetilde{{\mathbf{U}}}_{\beta,M}={\mathbf{U}}_{\beta,M}$ since $\Pi_{\beta}$ is bijective. ∎
We emphasize that when $M\geq 2$ the critical bases $\beta_{G}(M)$ and $\beta_{c}(M)$ defined in Definitions 1.1 and 1.2 are both strictly smaller than $\frac{M+2}{2}$ (see Figure 2 for $M\leq 10$). So, Proposition 2.5 implies that the critical bases for the phase transitions of $\widetilde{U}_{\beta,M}$ and $U_{\beta,M}$ are the same. Hence, in the following we only prove Theorem 1.4 for $\widetilde{U}_{\beta,M}$.
3. Generalized golden ratio $\beta_{G}(M)$ for a general $M$
Recall from Definition 1.1 the generalized golden ratio $\beta_{G}(M)$. Then by Lemma 2.1 it follows that the quasi-greedy $\beta_{G}(M)$-expansion of $1$ is given by
(3.1) $\delta(\beta_{G}(M))=\left\{\begin{array}[]{lll}((N+1)NN)^{\infty}&\textrm{if}&M=3N+1,\\ ((N+1)(N+1)N)^{\infty}&\textrm{if}&M=3N+2,\\ (N+1)^{\infty}&\textrm{if}&M=3N+3.\end{array}\right.$
The following result shows that $\beta_{G}(M)$ is the first critical base for $\widetilde{{\mathbf{U}}}_{\beta,M}$, i.e., $\widetilde{{\mathbf{U}}}_{\beta,M}$ is finite if and only if $\beta\leq\beta_{G}(M)$. When $M=1$, this was proven by Sidorov [28, Theorem 4.1] (see also, [23]).
Proposition 3.1.
Let $\beta_{G}(M)$ be defined as in Definition 1.1. Then $\beta_{G}(M)$ is a Perron number, and
#$\#\widetilde{{\mathbf{U}}}_{\beta,M}<+\infty\quad\Longleftrightarrow\quad\beta\leq\beta_{G}(M).$
Furthermore, for $\beta\in(1,\beta_{G}(M)]$ we have
$\widetilde{{\mathbf{U}}}_{\beta,M}=\bigcup_{j=0}^{2}\left\{\theta^{j}(\alpha_{00}^{\infty})\right\}=\left\{\alpha_{00}^{\infty},\alpha_{0M}^{\infty},\alpha_{M0}^{\infty}\right\}.$
Proof.
By (3.1) it follows that $\beta_{G}(M)$ is a Perron number. For the remaining statements we consider three cases: (I) $M=3N+1$; (II) $M=3N+2$; and (III) $M=3N+3$ for some $N\in\mathbb{N}_{0}$.
Case (I). $M=3N+1$ with $N\in\mathbb{N}_{0}$. Then by (3.1), Lemma 2.1 and Proposition 2.3 we only need to prove that $\widetilde{{\mathbf{U}}}_{\beta,M}=\left\{\alpha_{00}^{\infty},\alpha_{M0}^{\infty},\alpha_{0M}^{\infty}\right\}$ if $\delta(\beta)\preccurlyeq((N+1)NN)^{\infty}$; and #$\#\widetilde{{\mathbf{U}}}_{\beta,M}=\infty$ if $\delta(\beta)\succ((N+1)NN)^{\infty}$.
Take $\beta\in(1,M+1]$ so that $\delta(\beta)\preccurlyeq((N+1)NN)^{\infty}$. Let $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta,M}$, and suppose $d_{1}^{1},d_{1}^{2},\overline{d_{1}^{\oplus}}<M$. Then by Proposition 2.3 and (2.1) it follows that
(3.2) $d_{2}\in\left\{\alpha_{(N+1)(N+1)},\alpha_{(N-1)(N+1)},\alpha_{(N+1)(N-1)}\right\}=\left\{\theta^{j}(\alpha_{(N+1)(N+1)}):j=0,1,2\right\}$
or
(3.3) $d_{2}\in\left\{\alpha_{(N+1)N},\alpha_{N(N+1)},\alpha_{NN}\right\}=\left\{\theta^{j}(\alpha_{(N+1)N}):j=0,1,2\right\}.$
For (3.2) we assume by (2.7) that $d_{2}=\alpha_{(N+1)(N+1)}$. Then by Proposition 2.3 and (2.1) we obtain that
$d_{2}d_{3}d_{4}=\alpha_{(N+1)(N+1)}\alpha_{NN}\alpha_{NN},$
which yields $\overline{d_{2}^{\oplus}d_{3}^{\oplus}d_{4}^{\oplus}}=(N-1)(N+1)(N+1)$. This leads to a contradiction with $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta,M}$. For (3.3) we assume by (2.7) that $d_{2}=\alpha_{(N+1)N}$. Then by Proposition 2.3 and (2.1) it follows that
$d_{3}\in\left\{\alpha_{N(N+1)},\alpha_{NN},\alpha_{(N-1)(N+1)}\right\}.$
If $d_{3}=\alpha_{(N-1)(N+1)}$, then by a similar argument as in (3.2) we will obtain that $d_{3}d_{4}d_{5}=\alpha_{(N-1)(N+1)}\alpha_{(N+1)N}\alpha_{(N+1)N}$, leading to a contradiction with $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta,M}$. So, $d_{2}d_{3}\in\left\{\alpha_{(N+1)N}\alpha_{N(N+1)},\alpha_{(N+1)N}\alpha_{NN}\right\}$. By Proposition 2.3 and the same argument as above we can deduce that
$\left(\begin{array}[]{c}d_{2}^{1}d_{3}^{1}\ldots\\ d_{2}^{2}d_{3}^{2}\ldots\\ \overline{d_{2}^{\oplus}}\overline{d_{3}^{\oplus}}\ldots\end{array}\right)\in\left\{\left(\begin{array}[]{c}((N+1)NN)^{\infty}\\ (N(N+1)N)^{\infty}\\ (NN(N+1))^{\infty}\end{array}\right),\left(\begin{array}[]{c}((N+1)NN)^{\infty}\\ (NN(N+1))^{\infty}\\ (N(N+1)N)^{\infty}\end{array}\right)\right\},$
which again leads to a contradiction with $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta,M}$.
Therefore, $d_{1}^{1}=M,d_{1}^{2}=M$ or $\overline{d_{1}^{\oplus}}=M$. By (2.7) we may assume $d_{1}^{1}=M$. Then by (2.1) we have $d_{1}^{2}=\overline{d_{1}^{\oplus}}=0$. By Proposition 2.3 and the same argument as above we can prove that $(d_{i})=\alpha_{M0}^{\infty}$. Therefore, by (2.7) this proves $\widetilde{{\mathbf{U}}}_{\beta,M}=\left\{\alpha_{00}^{\infty},\alpha_{0M}^{\infty},\alpha_{M0}^{\infty}\right\}$.
On the other hand, if $\beta\in(1,M+1]$ with $\delta(\beta)\succ((N+1)NN)^{\infty}$ then we can prove that
$\alpha_{M0}^{k}(\alpha_{(N+1)N}\alpha_{N(N+1)}\alpha_{NN})^{\infty}\in\widetilde{{\mathbf{U}}}_{\beta,M}\quad\textrm{for any }k\in\mathbb{N}_{0}.$
So #$\#\widetilde{{\mathbf{U}}}_{\beta,M}=\infty$.
Case (II). $M=3N+2$ with $N\in\mathbb{N}_{0}$. By (3.1), Lemma 2.1 and Proposition 2.3 it suffices to prove that $\widetilde{{\mathbf{U}}}_{\beta,M}=\left\{\alpha_{00}^{\infty},\alpha_{M0}^{\infty},\alpha_{0M}^{\infty}\right\}$ if $\delta(\beta)\preccurlyeq((N+1)(N+1)N)^{\infty}$; and #$\#\widetilde{{\mathbf{U}}}_{\beta,M}=\infty$ if $\delta(\beta)\succ((N+1)(N+1)N)^{\infty}$.
Take $\beta\in(1,M+1]$ with $\delta(\beta)\preccurlyeq((N+1)(N+1)N)^{\infty}$, and let $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta,M}$. If $d_{1}^{1},d_{1}^{2},\overline{d_{1}^{\oplus}}<M$, then by Proposition 2.3 and (2.1) it follows that
$d_{2}\in\left\{\alpha_{(N+1)(N+1)},\alpha_{(N+1)N},\alpha_{N(N+1)}\right\}=\left\{\theta^{j}(\alpha_{(N+1)(N+1)}):j=0,1,2\right\}.$
Then by (2.7) we may assume $d_{2}=\alpha_{(N+1)(N+1)}$. By Proposition 2.3 and (2.1) we obtain that
$d_{2}d_{3}\ldots=(\alpha_{(N+1)(N+1)}\alpha_{(N+1)N}\alpha_{N(N+1)})^{\infty}\quad\textrm{or}\quad(\alpha_{(N+1)(N+1)}\alpha_{N(N+1)}\alpha_{(N+1)N})^{\infty}.$
Both cases will lead to a contradiction with $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta,M}$. Similar to Case $M=3N+1$, we can deduce that $\widetilde{{\mathbf{U}}}_{\beta,M}=\left\{\alpha_{00}^{\infty},\alpha_{M0}^{\infty},\alpha_{0M}^{\infty}\right\}$. On the other hand, if $\beta\in(1,M+1]$ with $\delta(\beta)\succ((N+1)(N+1)N)^{\infty}$, then we can prove that
$\alpha_{M0}^{k}(\alpha_{(N+1)(N+1)}\alpha_{N(N+1)}\alpha_{(N+1)N})^{\infty}\in\widetilde{{\mathbf{U}}}_{\beta,M}\quad\textrm{for any }k\in\mathbb{N}_{0}.$
So #$\#\widetilde{{\mathbf{U}}}_{\beta,M}=\infty$.
Case (III). $M=3N+3$ with $N\in\mathbb{N}_{0}$. By (3.1), Lemma 2.1 and Proposition 2.3 it suffices to prove that $\widetilde{{\mathbf{U}}}_{\beta,M}=\left\{\alpha_{00}^{\infty},\alpha_{0M}^{\infty},\alpha_{M0}^{\infty}\right\}$ if $\delta(\beta)\preccurlyeq(N+1)^{\infty}$, and #$\#\widetilde{{\mathbf{U}}}_{\beta,M}=\infty$ if $\delta(\beta)\succ(N+1)^{\infty}$.
Take $\beta\in(1,M+1]$ such that $\delta(\beta)\preccurlyeq(N+1)^{\infty}$, and take $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta,M}$. Note that $d_{1}\in\Omega_{M}$. If $d_{1}^{1},d_{1}^{2},\overline{d_{1}^{\oplus}}<M=3(N+1)$, then by Proposition 2.3 and (2.1) it follows that $d_{2}^{1}=d_{2}^{2}=\overline{d_{2}^{\oplus}}=N+1$. Since $N+1<M$, by the same argument we can deduce that
$d_{2}^{1}d_{3}^{1}\ldots=d_{2}^{2}d_{3}^{2}\ldots=\overline{d_{2}^{\oplus}}\overline{d_{3}^{\oplus}}\ldots=(N+1)^{\infty}\succcurlyeq\delta(\beta),$
leading to a contradiction with $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta,M}$. So, $d_{1}^{1}=M,d_{1}^{2}=M$ or $\overline{d_{1}^{\oplus}}=M$. By the same argument as that for $M=3N+1$ we can prove $\widetilde{{\mathbf{U}}}_{\beta,M}=\left\{\alpha_{00}^{\infty},\alpha_{0M}^{\infty},\alpha_{M0}^{\infty}\right\}$. On the other hand, take $\beta\in(1,M+1]$ so that $\delta(\beta)\succ(N+1)^{\infty}$. Then one can verify by Proposition 2.3 that
$\alpha_{M0}^{k}\alpha_{(N+1)(N+1)}^{\infty}\in\widetilde{{\mathbf{U}}}_{\beta,M}\quad\forall k\in\mathbb{N}_{0},$
which implies #$\#\widetilde{{\mathbf{U}}}_{\beta,M}=\infty$. This completes the proof. ∎
4. Generalized Komornik-Loreti constant $\beta_{c}(M)$ for $M=3N+1$
Let $M=3N+1$ with $N\in\mathbb{N}_{0}$. Recall from Definition 1.2 the generalized Komornik-Loreti constant $\beta_{c}(M)$. In this section we will show that $\dim_{H}\widetilde{{\mathbf{U}}}_{\beta,M}>0$ if and only if $\beta>\beta_{c}(M)$; and prove Theorem 1.4 for $M=3N+1$. When $M=1$, Theorem 1.4 was established by Kong and Li [23]. In this section we may assume $M=3N+1$ with $N\in\mathbb{N}$. The new challenge is that for $M=3N+1$ with $N\geq 1$ the digit $N-1$ might occur in $d_{n}^{1},d_{n}^{2}$ or $\overline{d_{n}^{\oplus}}$ for a sequence $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{c},M}$, while for $M=1$ this can never happen. Fortunately, we show in Lemma 4.5 that this digit $N-1$ can occur at most 3 times in any sequence of $\widetilde{{\mathbf{U}}}_{\beta_{c},M}$.
4.1. Properties of $\beta_{c}(M)$
Recall that $(\tau_{i})_{i=0}^{\infty}=0110100110010110\ldots\in\left\{0,1\right\}^{\mathbb{N}}$ is the classical Thue-Morse sequence (cf. [2]). Then $\tau_{2i+1}=1-\tau_{i}$ and $\tau_{2i}=\tau_{i}$ for all $i\geq 0$. Now we introduce a Thue-Morse type sequence $(\lambda_{i})\in\left\{N,N+1\right\}^{\mathbb{N}}$ such that
(4.1) $\lambda_{3k+1}=\tau_{2k+1}+N,\quad\lambda_{3k+2}=N,\quad\lambda_{3k+3}=\tau_{2k+2}+N$
for all $k\geq 0$. Then $(\lambda_{i})$ begins with $(N+1)N(N+1)NN(N+1)NNN(N+1)N(N+1)$.
The Thue-Morse type sequence $(\lambda_{i})$ can also be obtained by substitutions.
Lemma 4.1.
Let $(\tau_{i})_{i=0}^{\infty}$ be the classical Thue-Morse sequence. Then
$(\lambda_{i})_{i=1}^{\infty}=\sigma(\varphi_{1}(\tau_{0})\varphi_{1}(\tau_{1})\varphi_{1}(\tau_{2})\ldots),$
[…]
Proof.
Set $\lambda_{0}=N$. Then $\lambda_{0}=N+\tau_{0}$. By (4.1) and the definition of $(\tau_{i})$ it follows that
(4.2) $\lambda_{3k}=\tau_{k}+N,\quad\lambda_{3k+1}=1-\tau_{k}+N,\quad\lambda_{3k+2}=N$
[…]
$(\tau_{i})_{i=0}^{\infty}=\phi(\tau_{0})\phi(\tau_{1})\phi(\tau_{2})\ldots.$
[…]
$(\lambda_{i})_{i=0}^{\infty}=\varphi_{1}(\tau_{0})\varphi_{1}(\tau_{1})\varphi_{1}(\tau_{2})\ldots,$
completing the proof. ∎
Next we show that $(\lambda_{i})$ is the quasi-greedy expansion of $1$ in base $\beta_{c}(M)$.
Proposition 4.2.
Let $M=3N+1$. The quasi-greedy $\beta_{c}(M)$-expansion of $1$ is given by
$\delta(\beta_{c}(M))=\lambda_{1}\lambda_{2}\lambda_{3}\ldots=(N+1)N(N+1)NN(N+1)NNN\ldots,$
where $(\lambda_{i})$ is defined in (4.1). Furthermore, $\beta_{c}(M)$ is a transcendental number.
Let $(\overline{\tau_{i}})_{i=0}^{\infty}=(1-\tau_{i})=10010110\ldots\in\left\{0,1\right\}^{\mathbb{N}}$ be the reflection of $(\tau_{i})$. Set
(4.3) ${\Sigma}_{1}:=\left\{(N+1)N(N+1),NNN,(N+1)NN,NN(N+1)\right\}.$
We define a block map $\phi_{1}:{\Sigma}_{1}\rightarrow{\Sigma}_{1}$ by
(4.4) $\begin{split}(N+1)N(N+1)\quad&\longleftrightarrow\quad NNN,\qquad(N+1)NN\quad\longleftrightarrow\quad NN(N+1).\end{split}$
Then for a sequence $\mathbf{a}=a_{1}a_{2}\ldots\in{\Sigma}_{1}^{\mathbb{N}}$ we set $\phi_{1}(\mathbf{a}):=\phi_{1}(a_{1})\phi_{1}(a_{2})\cdots\in\Sigma_{1}^{\mathbb{N}}$. We emphasize that each digit $a_{i}$ is a block of length $3$ from ${\Sigma}_{1}$. Note that $(\lambda_{i})\in\Sigma_{1}^{\mathbb{N}}$. We define its conjugacy by
$(\hat{\lambda}_{i})_{i=1}^{\infty}:=\phi_{1}(\lambda_{1}\lambda_{2}\lambda_{3})\phi_{1}(\lambda_{4}\lambda_{5}\lambda_{6})\ldots\phi_{1}(\lambda_{3k+1}\lambda_{3k+2}\lambda_{3k+3})\ldots\in\Sigma_{1}^{\mathbb{N}}.$
Then $(\hat{\lambda}_{i})$ begins with $NNN(N+1)NN(N+1)N(N+1)NNN$. Comparing the definitions of $(\overline{\tau_{i}})$ and $(\hat{\lambda}_{i})$ we obtain that for any $k\in\mathbb{N}_{0}$,
(4.5) $\hat{\lambda}_{3k+1}=\overline{\tau_{2k+1}}+N,\quad\hat{\lambda}_{3k+2}=N,\quad\hat{\lambda}_{3k+3}=\overline{\tau_{2k+2}}+N.$
Lemma 4.3.
For any $n\in\mathbb{N}_{0}$ and any $0\leq i<3\cdot 2^{n}$, we have
(4.6) $\hat{\lambda}_{1}\ldots\hat{\lambda}_{3\cdot 2^{n}-i}$ $\prec\lambda_{i+1}\ldots\lambda_{3\cdot 2^{n}}\preccurlyeq\lambda_{1}\ldots\lambda_{3\cdot 2^{n}-i},$
(4.7) $\hat{\lambda}_{1}\ldots\hat{\lambda}_{3\cdot 2^{n}-i}$ $\preccurlyeq\hat{\lambda}_{i+1}\ldots\hat{\lambda}_{3\cdot 2^{n}}\prec\lambda_{1}\ldots\lambda_{3\cdot 2^{n}-i}.$
Proof.
Since the proof of (4.7) is similar to (4.6), we only prove (4.6). Take $n\in\mathbb{N}_{0}$ and $i\in\left\{0,1,\ldots,3\cdot 2^{n}-1\right\}$. We will prove (4.6) by the following three cases: (I) $i=3k$, (II) $i=3k+1$, and (III) $i=3k+2$ for some $0\leq k<2^{n}$.
Case (I). $i=3k$ with $0\leq k<2^{n}$. Then by (4.1) we have
(4.8) $\lambda_{i+1}\ldots\lambda_{3\cdot 2^{n}}=(\tau_{2k+1}+N)N(\tau_{2k+2}+N)\cdots(\tau_{2^{n+1}-1}+N)N(\tau_{2^{n+1}}+N).$
Note by [22] that
(4.9) $\overline{\tau_{1}\ldots\tau_{2^{n+1}-j}}\prec\tau_{j+1}\ldots\tau_{2^{n+1}}\preccurlyeq\tau_{1}\ldots\tau_{2^{n+1}-j}\quad\forall 0\leq j<2^{n+1}.$
Then by (4.8), (4.9), (4.1) and (4.5) it follows that
$\lambda_{i+1}\ldots\lambda_{3\cdot 2^{n}}\preccurlyeq(\tau_{1}+N)N(\tau_{2}+N)\cdots(\tau_{2^{n+1}-2k-1}+N)N(\tau_{2^{n+1}-2k}+N)=\lambda_{1}\ldots\lambda_{3\cdot 2^{n}-i}$
and
$\lambda_{i+1}\ldots\lambda_{3\cdot 2^{n}}\succ(\overline{\tau_{1}}+N)N(\overline{\tau_{2}}+N)\cdots(\overline{\tau_{2^{n+1}-2k-1}}+N)N(\overline{\tau_{2^{n+1}-2k}}+N)=\hat{\lambda}_{1}\ldots\hat{\lambda}_{3\cdot 2^{n}-i}$
as desired.
Case (II). $i=3k+1$ with $0\leq k<2^{n}$. If $k=2^{n}-1$, then
$\lambda_{i+1}\lambda_{i+2}=\lambda_{3\cdot 2^{n}-1}\lambda_{3\cdot 2^{n}}=N(\tau_{2^{n+1}}+N)=N(N+1),$
where the last equality follows by using $\tau_{2^{m}}=1$ for all $m\geq 0$. Note that $\lambda_{1}\lambda_{2}=(N+1)N$ and $\hat{\lambda}_{1}\hat{\lambda}_{2}=NN$. Then it is easy to verify (4.6) when $i=3\cdot(2^{n}-1)+1$.
Next we assume $10\leq k<2^{n}-$ 1. Then by (4.1) we have $\lambda_{i+1}=\lambda_{3k+2}=N<N+1=\lambda_{1}$, which yields $\lambda_{i+1}\ldots\lambda_{3\cdot 2^{n}}\prec\lambda_{1}\ldots\lambda_{3\cdot 2^{n}-i}$. On the other hand, observe by (4.1) that
$\lambda_{i+1}\lambda_{i+2}\lambda_{i+3}$ $=N(\tau_{2k+2}+N)(\tau_{2k+3}+N)$
$=N(\tau_{k+1}+N)(1-\tau_{k+1}+N)$
$\succcurlyeq NN(N+1)\succ NNN=\hat{\lambda}_{1}\hat{\lambda}_{2}\hat{\lambda}_{3},$
where the second equality holds by the definition of $(\tau_{i})$ that $\tau_{2i}=\tau_{i}$ and $\tau_{2i+1}=1-\tau_{i}$ for all $i\geq 0$. Thus, $\lambda_{i+1}\ldots\lambda_{3\cdot 2^{n}}\succ\hat{\lambda}_{1}\ldots\hat{\lambda}_{3\cdot 2^{n}-i}$ as required.
Case (III). $i=3k+2$ with $0\leq k<2^{n}$. If $k=2^{n}-1$, then $\lambda_{i+1}=\lambda_{3\cdot 2^{n}}=N+1$. Since $\lambda_{1}=N+1$ and $\hat{\lambda}_{1}=N$, it is clear that (4.6) holds for $i=3\cdot(2^{n}-1)+2$. In the following we assume $10\leq k<2^{n}-$ 1. Then by (4.1) we obtain
$\lambda_{i+1}\lambda_{i+2}\lambda_{i+3}$ $=(\tau_{2k+2}+N)(\tau_{2k+3}+N)N$
$=(\tau_{k+1}+N)(1-\tau_{k+1}+N)N$
$\preccurlyeq(N+1)NN\prec(N+1)N(N+1)=\lambda_{1}\lambda_{2}\lambda_{3},$
and
$\lambda_{i+1}\lambda_{i+2}\lambda_{i+3}=(\tau_{k+1}+N)(1-\tau_{k+1}+N)N\succcurlyeq N(N+1)N\succ NNN=\hat{\lambda}_{1}\hat{\lambda}_{2}\hat{\lambda}_{3}.$
This proves (4.6), completing the proof. ∎
To prove Proposition 4.2 we also need the following result due to Mahler [25].
Lemma 4.4.
If $z$ is an algebraic number in the open unit disc, then
$Z=\sum_{i=1}^{\infty}\tau_{i}z^{i}$
is transcendental, where $(\tau_{i})$ is the classical Thue-Morse sequence.
Proof of Proposition 4.2.
Note by Definition 1.2 and (4.1) that $(\lambda_{i})$ is an expansion of $1$ in base $\beta_{c}=\beta_{c}(M)$. Then by Lemmas 2.1 and 4.3 it follows that $\delta(\beta_{c})=(\lambda_{i})$. So, by (4.1) we obtain
$1=\sum_{i=1}^{\infty}\frac{\lambda_{i}}{\beta_{c}^{i}}$ $=\sum_{k=0}^{\infty}\frac{\tau_{2k+1}+N}{\beta_{c}^{3k+1}}+\sum_{k=0}^{\infty}\frac{N}{\beta_{c}^{3k+2}}+\sum_{k=0}^{\infty}\frac{\tau_{2k+2}+N}{\beta_{c}^{3k+3}}$
$=\sum_{i=1}^{\infty}\frac{N}{\beta_{c}^{i}}+\sum_{k=0}^{\infty}\frac{1-\tau_{k}}{\beta_{c}^{3k+1}}+\sum_{k=0}^{\infty}\frac{\tau_{k+1}}{\beta_{c}^{3k+3}}$
$=\frac{N}{\beta_{c}-1}+\frac{\beta_{c}^{2}}{\beta_{c}^{3}-1}-\frac{1}{\beta_{c}}\sum_{k=1}^{\infty}\frac{\tau_{k}}{\beta_{c}^{3k}}+\sum_{k=1}^{\infty}\frac{\tau_{k}}{\beta_{c}^{3k}},$
where in the last equality we used $\tau_{0}=0$. Rearranging the above equation yields that
(4.10) $\sum_{k=1}^{\infty}\frac{\tau_{k}}{\beta_{c}^{3k}}=\frac{1-\frac{N}{\beta_{c}-1}-\frac{\beta_{c}^{2}}{\beta_{c}^{3}-1}}{1-\frac{1}{\beta_{c}}}.$
Suppose on the contrary that $\beta_{c}\in(N+1,N+2)$ is algebraic. Then so is $\beta_{c}^{-3}\in(0,1)$. By Lemma 4.4 it follows that the left-hand side of (4.10) is transcendental, while the right hand-side of (4.10) is algebraic, leading to a contradiction. So, $\beta_{c}$ is transcendental. ∎
4.2. Patterns in $\widetilde{{\mathbf{U}}}_{\beta_{c},M}$
We will construct a sequence of bases $(\beta_{n})_{n=0}^{\infty}$ strictly increasing to $\beta_{c}=\beta_{c}(M)$ such that the difference set $\widetilde{{\mathbf{U}}}_{\beta_{n+1},M}\setminus\widetilde{{\mathbf{U}}}_{\beta_{n},M}$ is countably infinite for all $n\geq 0$. In fact we will show that $\widetilde{{\mathbf{U}}}_{\beta_{n+1},M}\setminus\widetilde{{\mathbf{U}}}_{\beta_{n},M}$ consists of only eventually periodic sequences.
First we construct a sequence of blocks $({\mathbf{t}}_{n})$ approximating $(\lambda_{i})=\delta(\beta_{c})$ componentwisely. Recall from (4.3) that $\Sigma_{1}=\left\{(N+1)N(N+1),NNN,(N+1)NN,NN(N+1)\right\}$. Let
(4.11) ${\mathbf{t}}_{0}=(N+1)NN,\quad\textrm{and}\quad{\mathbf{t}}_{n+1}:={\mathbf{t}}_{n}^{+}\phi_{1}({\mathbf{t}}_{n}^{+})\quad\forall n\geq 0,$
where $\phi_{1}$ is the block map defined in (4.4), and for a word $\omega=\omega_{1}\ldots\omega_{m}\in\left\{N,N+1\right\}^{m}$ we set $\omega^{+}=\omega_{1}\ldots\omega_{m-1}(\omega_{m}+1)$. Then ${\mathbf{t}}_{n}\in\Sigma_{1}^{2^{n}}$ for any $n\geq 0$. For example,
${\mathbf{t}}_{1}=(N+1)N(N+1)NNN,\quad{\mathbf{t}}_{2}=(N+1)N(N+1)NN(N+1)\,NNN(N+1)NN.$
Observe that each ${\mathbf{t}}_{n}^{+}$ is a prefix of $(\lambda_{i})$, i.e., ${\mathbf{t}}_{n}^{+}=\lambda_{1}\ldots\lambda_{3\cdot 2^{n}}$. By Lemma 4.3 it follows that $\sigma^{i}({\mathbf{t}}_{n}^{\infty})\preccurlyeq{\mathbf{t}}_{n}^{\infty}$ for all $i\geq 0.$ So, by Lemma 2.1 there exists a unique $\beta_{n}=\beta_{n}(M)\in(1,M+1]$ such that
(4.12) $\delta(\beta_{n})={\mathbf{t}}_{n}^{\infty}.$
Note by Proposition 3.1 that $\delta(\beta_{G}(M))=((N+1)NN)^{\infty}={\mathbf{t}}_{0}^{\infty}=\delta(\beta_{0}(M)).$ Then $\beta_{G}(M)=\beta_{0}(M)$ by Lemma 2.1. Observe that $\delta(\beta_{n})={\mathbf{t}}_{n}^{\infty}\nearrow(\lambda_{i})=\delta(\beta_{c})$ as $n\to\infty$. By Lemma 2.1 it follows that
$\beta_{G}=\beta_{0}<\beta_{1}<\cdots<\beta_{n}<\beta_{n+1}<\cdots,\quad\textrm{and}\quad\beta_{n}\nearrow\beta_{c}\quad\textrm{as }n\to\infty.$
In the following we describe the block patterns appearing in sequences of $\widetilde{{\mathbf{U}}}_{\beta_{c},M}$. Let
$\pazocal A_{1}$ :=$:=\left\{\alpha_{(N+1)N}\alpha_{N(N+1)}\alpha_{(N+1)N},\quad\alpha_{NN}\alpha_{N(N+1)}\alpha_{NN},\right.$
$\hskip 39.83368pt\left.\alpha_{(N+1)N}\alpha_{N(N+1)}\alpha_{NN},\quad\alpha_{NN}\alpha_{N(N+1)}\alpha_{(N+1)N}\right\}.$
We define the substitution $\Phi_{1}:\pazocal A_{1}\to\pazocal A_{1}$ by
(4.13) $\begin{split}\alpha_{(N+1)N}\alpha_{N(N+1)}\alpha_{(N+1)N}\quad&\longleftrightarrow\quad\alpha_{NN}\alpha_{N(N+1)}\alpha_{NN},\\ \alpha_{(N+1)N}\alpha_{N(N+1)}\alpha_{NN}\quad&\longleftrightarrow\quad\alpha_{NN}\alpha_{N(N+1)}\alpha_{(N+1)N}.\end{split}$
By using $\Phi_{1}$ we construct a sequence of blocks $\left\{\mathbf{a}_{n}\right\}_{n=0}^{\infty}$ as follows:
(4.14) $\mathbf{a}_{0}=\alpha_{(N+1)N}\alpha_{N(N+1)}\alpha_{NN},\quad\textrm{and}\quad\mathbf{a}_{n+1}=\mathbf{a}_{n}^{+}\Phi_{1}(\mathbf{a}_{n}^{+})\quad\textrm{for }n\geq 0,$
where $\mathbf{a}_{n}^{+}$ is the word in $\pazocal A_{1}^{*}$ by changing the last digit of $\mathbf{a}_{n}$ from $\alpha_{NN}$ to $\alpha_{(N+1)N}$. Then by the definition of $\Phi_{1}$ each $\mathbf{a}_{n}$ has length $3\cdot 2^{n}$ and ends with $\alpha_{NN}$. If we write $\mathbf{a}_{n}=a_{1}\ldots a_{3\cdot 2^{n}}$, then $a_{3k+2}=a_{N(N+1)}$ for all $0\leq k<2^{n}$. By (4.14) it follows that
$\mathbf{a}_{0}\sim\left(\begin{array}[]{ccc}N+1&N&N\\ N&N+1&N\\ N&N&N+1\end{array}\right),\quad\mathbf{a}_{1}\sim\left(\begin{array}[]{cc}(N+1)N(N+1)&NNN\\ N(N+1)N&N(N+1)N\\ NNN&(N+1)N(N+1)\end{array}\right),$
and
$\mathbf{a}_{2}\sim\left(\begin{array}[]{cccc}(N+1)N(N+1)&NN(N+1)&NNN&(N+1)NN\\ N(N+1)N&N(N+1)N&N(N+1)N&N(N+1)N\\ NNN&(N+1)NN&(N+1)N(N+1)&NN(N+1)\end{array}\right).$
Comparing the definitions of $({\mathbf{t}}_{n})$ and $(\mathbf{a}_{n})$ in (4.11) and (4.14) respectively, it follows that
(4.15) $\mathbf{a}_{n}^{1}={\mathbf{t}}_{n},\quad\mathbf{a}_{n}^{2}=(N(N+1)N)^{n}\quad\textrm{and}\quad\overline{\mathbf{a}_{n}^{\oplus}}=\phi_{1}({\mathbf{t}}_{n})$
for any $n\geq 0$.
The big difference between $M=1$ and $M=3N+1$ with $N\geq 1$ is that in the later case the digit $N-1$ might occur in $d_{n}^{1}=N-1,d_{n}^{2}=N-1$ or $\overline{d_{n}^{\oplus}}=N-1$ for some sequence $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{c},M}$. In other words, when $M=3N+1$ with $N\geq 1$, sequences in $\widetilde{{\mathbf{U}}}_{\beta_{c},M}$ might contain digit from
(4.16) $\mathscr{E}_{1}:=\bigcup_{j=0}^{2}\left\{\theta^{j}(\alpha_{(N+1)(N+1)})\right\}=\left\{\alpha_{(N+1)(N+1)},\alpha_{(N-1)(N+1)},\alpha_{(N+1)(N-1)}\right\}.$
Fortunately, we can prove that any sequence of $\widetilde{{\mathbf{U}}}_{\beta_{c},M}$ contains at most three digits from $\mathscr{E}_{1}$.
Lemma 4.5.
Let $M=3N+1$ with $N\geq 1$. If $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{c},M}$, then
#$\#\left\{i\in\mathbb{N}:d_{i}\in\mathscr{E}_{1}\right\}\leq 3.$
Proof.
Take $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{c},M}$. Note by Proposition 3.1 that $\widetilde{{\mathbf{U}}}_{\beta_{c},M}\supset\widetilde{{\mathbf{U}}}_{\beta_{0},M}=\left\{\alpha_{00}^{\infty},\alpha_{0M}^{\infty},\alpha_{M0}^{\infty}\right\}.$ If $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{0},M}$, then we are done. So in the following we assume $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{c},M}\setminus\widetilde{{\mathbf{U}}}_{\beta_{0},M}$. Note by Proposition 4.2 that $\delta(\beta_{c})\prec(N+1)^{\infty}$. By Proposition 2.3 and (2.1) there exists a smallest $m\in\mathbb{N}$ such that
(4.17) $d_{m+k}^{1},\leavevmode\nobreak\ d_{m+k}^{2},\leavevmode\nobreak\ \overline{d_{m+k}^{\oplus}}\leq N+1<M\quad\forall k\geq 0.$
Suppose on the contrary there exist $0\leq p_{1}<p_{2}<p_{3}<p_{4}$ such that
(4.18) $d_{m+p_{j}}\in\left\{\alpha_{(N+1)(N+1)},\alpha_{(N-1)(N+1)},\alpha_{(N+1)(N-1)}\right\}\quad\textrm{for all }1\leq j\leq 4.$
Write $p:=p_{4}$, and without loss of generality we assume $d_{m+p}=\alpha_{(N+1)(N+1)}$. Then by (4.17), Propositions 2.3 and 4.2 it follows that
$d_{m+p-1}d_{m+p}d_{m+p+1}=\alpha_{NN}\alpha_{(N+1)(N+1)}\alpha_{NN}.$
For $d_{m+p+2}$ we have three possibilities: $d_{m+p+2}=\alpha_{(N+1)N},\alpha_{N(N+1)}$ or $\alpha_{(N+1)(N+1)}$.
Claim 1. $d_{m+p+2}\neq\alpha_{(N+1)(N+1)}$.
Suppose on the contrary that $d_{m+p+2}=\alpha_{(N+1)(N+1)}$. Then by using $d_{m+p}d_{m+p+1}d_{m+p+2}=\alpha_{(N+1)(N+1)}\alpha_{NN}\alpha_{(N+1)(N+1)}$, Propositions 2.3 and 4.2 we must have
$d_{m+p+3}d_{m+p+4}=\alpha_{NN}\alpha_{NN}.$
This implies that $\overline{d_{m+p+2}^{\oplus}}=N-1<M$ and $\overline{d_{m+p+3}^{\oplus}d_{m+p+4}^{\oplus}}=(N+1)(N+1)\succ\delta_{1}(\beta_{c})\delta_{2}(\beta_{c})$, leading to a contradiction with $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{c},M}$. So, Claim 1 holds.
By Claim 1 we have $d_{m+p+2}=\alpha_{(N+1)N}$ or $\alpha_{N(N+1)}$. By symmetry we may assume $d_{m+p+2}=\alpha_{(N+1)N}$. Then by (4.17), Propositions 2.3 and 4.2 it follows that
$d_{m+p-2}=\alpha_{N(N+1)}\quad\textrm{and}\quad d_{m+p-3}\in\left\{\alpha_{(N+1)N},\alpha_{NN},\alpha_{(N+1)(N-1)}\right\}.$
Claim 2. $d_{m+p-3}\neq\alpha_{(N+1)(N-1)}$.
Suppose on the contrary that $d_{m+p-3}=\alpha_{(N+1)(N-1)}$. Then
$d_{m+p-3}\ldots d_{m+p+2}\quad\sim\quad\left(\begin{array}[]{cccccc}N+1&N&N&N+1&N&N+1\\ N-1&N+1&N&N+1&N&N\\ N+1&N&N+1&N-1&N+1&N\end{array}\right).$
By the definition of $p_{i}$ in (4.18) it follows that $p=p_{4}\geq 5$. So, by (4.17) Propositions 2.3 and 4.2 it follows that $d_{m+p-4}=\alpha_{N(N+1)}$ and $d_{m+p-5}=\alpha_{(N+1)N}$ or $\alpha_{NN}$. However, if $d_{m+p-5}=\alpha_{(N+1)N}$, then by Proposition 4.2 we have
$d_{m+p-5}^{1}\ldots d_{m+p+2}^{1}=(N+1)N(N+1)NN(N+1)N(N+1)\succ\delta_{1}(\beta_{c})\ldots\delta_{8}(\beta_{c}),$
leading to a contradiction. If $d_{m+p-5}=\alpha_{NN}$, then $\overline{d_{m+p-5}^{\oplus}}\ldots\overline{d_{m+p+1}^{\oplus}}=(N+1)N(N+1)N(N+1)\succ\delta_{1}(\beta_{c})\ldots\delta_{5}(\beta_{c})$, again leading to a contradiction. This proves Claim 2.
So, by Claim 2 we have $d_{m+p-3}\in\left\{\alpha_{(N+1)N},\alpha_{NN}\right\}$. Without loss of generality we may assume $d_{m+p-3}=\alpha_{NN}$. Then
$d_{m+p-3}\ldots d_{m+p+2}\quad\sim\quad\left(\begin{array}[]{cccccc}N&N&N&N+1&N&N+1\\ N&N+1&N&N+1&N&N\\ N+1&N&N+1&N-1&N+1&N\end{array}\right),$
which implies $p=p_{4}\geq 6$ by (4.18). By (4.17), Propositions 2.3 and 4.2 we obtain that $d_{m+p-5}d_{m+p-4}=\alpha_{N(N+1)}\alpha_{(N+1)N}$ and $d_{m+p-6}\in\left\{\alpha_{(N+1)N},\alpha_{NN},\alpha_{(N+1)(N-1)}\right\}$. Similar to Claim 2 we can prove that $d_{m+p-6}\neq\alpha_{(N+1)(N-1)}$.
Proceeding this argument we can prove that
$d_{m+p-3i}d_{m+p-3i+1}d_{m+p-3i+2}\in\left\{\alpha_{(N+1)N}\alpha_{N(N+1)}\alpha_{NN},\alpha_{NN}\alpha_{N(N+1)}\alpha_{NN}\right\}$
for all $1\leq i\leq\lfloor\frac{p}{3}\rfloor$. This leads to a contradiction with (4.18), completing the proof. ∎
Lemma 4.6.
Let $M=3N+1$ with $N\in\mathbb{N}_{0}$. Then for any sequence $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{c},M}\setminus\widetilde{{\mathbf{U}}}_{\beta_{0},M}$ there exists $m_{1}\in\mathbb{N}$ such that
$d_{k}\in\left\{\alpha_{(N+1)N},\alpha_{NN},\alpha_{N(N+1)}\right\}\quad\forall k>m_{1}.$
Proof.
Take $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{c},M}\setminus\widetilde{{\mathbf{U}}}_{\beta_{0},M}$. If $M=3N+1$ with $N=0$, then it is clear that $d_{i}\in\left\{\alpha_{(N+1)N},\alpha_{NN},\alpha_{N(N+1)}\right\}$ for all $i\geq 1$. In the following we assume $N\geq 1$. Note by Proposition 3.1 that $\widetilde{{\mathbf{U}}}_{\beta_{0},M}=\widetilde{{\mathbf{U}}}_{\beta_{G},M}=\left\{\alpha_{0M}^{\infty},\alpha_{M0}^{\infty},\alpha_{00}^{\infty}\right\}$. Furthermore, by Proposition 4.2 and Lemma 2.1 that $\delta(\beta_{c})\prec(N+1)^{\infty}$. Then there exists $m_{0}\in\mathbb{N}$ such that $d_{m_{0}}^{1}<M,d_{m_{0}}^{1}<M$ and $\overline{d_{m_{0}}^{\oplus}}<M$. Since $M=3N+1$ with $N\geq 1$, by (2.1) and Proposition 2.3 it follows that
$d_{k}\in\bigcup_{j=0}^{2}\left\{\theta^{j}(\alpha_{(N+1)N}),\theta^{j}(\alpha_{(N+1)(N+1)})\right\}\quad\forall k>m_{0}.$
Hence, by Lemma 4.5 there exists $m_{1}>m_{0}$ such that
$d_{k}\in\bigcup_{j=0}^{2}\left\{\theta^{j}(\alpha_{(N+1)N})\right\}=\left\{\alpha_{(N+1)N},\alpha_{N(N+1)},\alpha_{NN}\right\}\quad\forall k>m_{1},$
completing the proof. ∎
Now we are ready to describe the patterns appearing in sequences of $\widetilde{{\mathbf{U}}}_{\beta_{c},M}$.
Proposition 4.7.
Let $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{c},M}$ and $n\in\mathbb{N}_{0}$. Then there exists $m_{1}\in\mathbb{N}$ such that for all $m>m_{1}$ the following statements hold.
  1. (i)
    If $d^{1}_{m}<M$ and $d_{m+1}\ldots d_{m+3\cdot 2^{n}}=\mathbf{a}_{n}^{+}$, then $d_{m+3\cdot 2^{n}+1}\ldots d_{m+3\cdot 2^{n+1}}\in\left\{\Phi_{1}(\mathbf{a}_{n}),\Phi_{1}(\mathbf{a}_{n}^{+})\right\}.$
  2. (ii)
    If $d^{1}_{m}>0$ and $d_{m+1}\ldots d_{m+3\cdot 2^{n}}=\Phi_{1}(\mathbf{a}_{n}^{+})$, then $d_{m+3\cdot 2^{n}+1}\ldots d_{m+3\cdot 2^{n+1}}\in\left\{\mathbf{a}_{n},\mathbf{a}_{n}^{+}\right\}.$
Remark 4.8.
  • Note by the definition of $(\mathbf{a}_{n})$ in (4.14). Then Proposition 4.7 implies that if $d_{m}^{1}<M$ and $d_{m+1}\ldots d_{m+3\cdot 2^{n}}=\mathbf{a}_{n}^{+}$, then $d_{m+1}\ldots d_{m+3\cdot 2^{n+1}}=\mathbf{a}_{n+1}$ or $\mathbf{a}_{n+1}^{+}$. Similarly, if $d_{m}^{1}>0$ and $d_{m+1}\ldots d_{m+3\cdot 2^{n}}=\Phi_{1}(\mathbf{a}_{n}^{+})$, then $d_{m+1}\ldots d_{m+3\cdot 2^{n+1}}=\Phi_{1}(\mathbf{a}_{n+1})$ or $\Phi_{1}(\mathbf{a}_{n+1}^{+})$.
  • Observe by (2.7) that $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta,M}$ if and only if $\theta((d_{i}))\in\widetilde{{\mathbf{U}}}_{\beta,M}$, which is also equivalent to $\theta^{2}((d_{i}))\in\widetilde{{\mathbf{U}}}_{\beta,M}$. So, Proposition 4.7 also holds if we replace $(d_{i})$ by $\theta((d_{i}))$ or $\theta^{2}((d_{i}))$. For example, if $\overline{d_{m}^{\oplus}}<M$ and $d_{m+1}\ldots d_{m+3\cdot 2^{n}}=\theta(\mathbf{a}_{n}^{+})$, then $d_{m+3\cdot 2^{n}+1}\ldots d_{m+3\cdot 2^{n+1}}=\theta(\Phi_{1}(\mathbf{a}_{n}))$ or $\theta(\Phi_{1}(\mathbf{a}_{n}^{+}))$.
First we prove Proposition 4.7 for $n=0$ and $n=1$.
Lemma 4.9.
Let $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{c},M}$. Then there exists $m_{1}\in\mathbb{N}$ such that for all $m>m_{1}$ the following statements hold.
  1. (i)
    If $d^{1}_{m}<M$ and $d_{m+1}d_{m+2}d_{m+3}=\mathbf{a}_{0}^{+}$, then $d_{m+4}d_{m+5}d_{m+6}=\Phi_{1}(\mathbf{a}_{0})$ or $\Phi_{1}(\mathbf{a}_{0}^{+}).$
  2. (ii)
    If $d^{1}_{m}>0$ and $d_{m+1}d_{m+2}d_{m+3}=\Phi_{1}(\mathbf{a}_{0}^{+})$, then $d_{m+4}d_{m+5}d_{m+6}=\mathbf{a}_{0}$ or $\mathbf{a}_{0}^{+}.$
Proof.
Since the proof of (ii) is similar, we only prove (i). Take $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{c},M}$. If $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{0},M}$, then by Proposition 3.1 there is nothing to prove. In the following we assume $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{c},M}\setminus\widetilde{{\mathbf{U}}}_{\beta_{0},M}$. Then by Lemma 4.6 there exists $m_{1}\in\mathbb{N}$ such that
(4.19) $d_{j}\in\left\{\alpha_{(N+1)N},\alpha_{N(N+1)},\alpha_{NN}\right\}\quad\forall j>m_{1}.$
Take $m>m_{1}$, and suppose $d_{m+1}d_{m+2}d_{m+3}=\mathbf{a}_{0}^{+}=\alpha_{(N+1)N}\alpha_{N(N+1)}\alpha_{(N+1)N}$ with $d_{m}^{1}<M$. Then it suffices to prove
(4.20) $d_{m+4}d_{m+5}d_{m+6}\in\left\{\alpha_{NN}\alpha_{N(N+1)}\alpha_{NN},\quad\alpha_{NN}\alpha_{N(N+1)}\alpha_{(N+1)N}\right\}.$
We first prove $d_{m+4}=\alpha_{NN}$. Suppose on the contrary that $d_{m+4}\neq\alpha_{NN}$. Note that $d_{m+1}d_{m+2}d_{m+3}=\mathbf{a}_{0}^{+}$ and $d_{m}^{1}<M$. Moreover, observe by Proposition 4.2 that $\delta(\beta_{c})$ begins with $(N+1)N(N+1)NN(N+1)NNN$. Then by (2.1), (4.19) and Proposition 2.3 it follows that
$d_{m+1}\ldots d_{m+9}\sim\left(\begin{array}[]{ccc}(N+1)N(N+1)&NN(N+1)&NNN\\ N(N+1)N&(N+1)NN&N(N+1)(N+1)\\ NNN&N(N+1)N&(N+1)NN\end{array}\right)$
or
$d_{m+1}\ldots d_{m+9}\sim\left(\begin{array}[]{ccc}(N+1)N(N+1)&NN(N+1)&NNN\\ N(N+1)N&(N+1)NN&(N+1)NN\\ NNN&N(N+1)N&N(N+1)(N+1)\end{array}\right).$
However, in the first case we have $d_{7}^{2}=N<M$ and $d_{8}^{2}d_{9}^{2}=(N+1)(N+1)\succ\delta_{1}(\beta_{c})\delta_{2}(\beta_{c})$; and in the second case we have $\overline{d_{7}^{\oplus}}=N<M$ and $\overline{d_{8}^{\oplus}d_{9}^{\oplus}}=(N+1)(N+1)\succ\delta_{1}(\beta_{c})\delta_{2}(\beta_{c})$. Both cases lead to a contradiction with $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{c},M}$. So, $d_{m+4}=\alpha_{NN}$.
Note that $d_{m+1}\ldots d_{m+4}=\alpha_{(N+1)N}\alpha_{N(N+1)}\alpha_{(N+1)N}\alpha_{NN}$. Then by (2.1), (4.14) and Proposition 2.3 it follows that $d_{m+5}d_{m+6}\in\left\{\alpha_{N(N+1)}\alpha_{NN},\alpha_{N(N+1)}\alpha_{(N+1)N}\right\}$, proving (4.20). ∎
Lemma 4.10.
Let $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{c},M}$. Then there exists $m_{1}\in\mathbb{N}$ such that for all $m>m_{1}$ the following statements hold.
  1. (i)
    If $d^{1}_{m}<M$ and $d_{m+1}\ldots d_{m+6}=\mathbf{a}_{1}^{+}$, then $d_{m+7}\ldots d_{m+12}=\Phi_{1}(\mathbf{a}_{1})$ or $\Phi_{1}(\mathbf{a}_{1}^{+}).$
  2. (ii)
    If $d^{1}_{m}>0$ and $d_{m+1}\ldots d_{m+6}=\Phi_{1}(\mathbf{a}_{1}^{+})$, then $d_{m+7}\ldots d_{m+12}=\mathbf{a}_{1}$ or $\mathbf{a}_{1}^{+}.$
Proof.
Since the proof of (ii) is similar, we only prove (i). Take $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{c},M}$. As in the proof of Lemma 4.9 we may assume $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{c},M}\setminus\widetilde{{\mathbf{U}}}_{\beta_{0},M}$. Then by Lemma 4.6 there exists $m_{1}\in\mathbb{N}$ such that $d_{j}\in\left\{\alpha_{(N+1)N},\alpha_{N(N+1)},\alpha_{NN}\right\}$ for all $j>m_{1}$. Take $m>m_{1}$, and suppose $d_{m+1}\ldots d_{m+6}=\mathbf{a}_{1}^{+}$ with $d_{m}^{1}<M$. Note by (4.15) that $d_{m+1}^{1}\ldots d_{m+6}^{1}={\mathbf{t}}_{1}^{+}=(N+1)N(N+1)NN(N+1)$. Then by Propositions 2.3 and 4.2 it follows that $d_{m+7}\in\left\{\alpha_{NN},\alpha_{N(N+1)}\right\}.$
Suppose $d_{m+7}=\alpha_{N(N+1)}$. Then by using $d_{m+1}\ldots d_{m+6}=\mathbf{a}_{1}^{+}$ and Proposition 4.2 it follows that
$d_{m+1}\ldots d_{m+9}\quad\sim\quad\left(\begin{array}[]{ccc}(N+1)N(N+1)&NN(N+1)&NNN\\ N(N+1)N&N(N+1)N&(N+1)NN\\ NNN&NNN&N(N+1)(N+1)\end{array}\right).$
This gives $\overline{d_{m+7}^{\oplus}}=N<M$ and $\overline{d_{m+8}^{\oplus}d_{m+9}^{\oplus}}=(N+1)(N+1)$, leading to a contradiction with $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{c},M}$. So, $d_{m+7}=\alpha_{NN}$.
Therefore, by Propositions 2.3 and 4.2 we can deduce that $d_{m+8}d_{m+9}=\alpha_{N(N+1)}\alpha_{NN}$, and thus $d_{m+7}d_{m+8}d_{m+9}=\alpha_{NN}\alpha_{N(N+1)}\alpha_{NN}=\Phi(\mathbf{a}_{0}^{+}).$ Since $d_{m+6}^{1}=N+1>0$, by Lemma 4.9 (ii) it follows that
$d_{m+7}\ldots d_{m+12}=\Phi_{1}(\mathbf{a}_{0}^{+})\mathbf{a}_{0}=\Phi_{1}(\mathbf{a}_{1}^{+})\quad\textrm{or}\quad\Phi_{1}(\mathbf{a}_{0}^{+})\mathbf{a}_{0}^{+}=\Phi_{1}(\mathbf{a}_{1}),$
completing the proof. ∎
Proof of Proposition 4.7.
Since the proof of (ii) is similar, we only prove (i). This will be done by induction on $n$. Note by Lemmas 4.9 and 4.10 that (i) holds for $n=0,1$. Now we assume (i) holds for all $k\leq n$ for some $n\in\mathbb{N}$, and we will prove it for $k=n+1$. Take $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{c},M}$. As in the proof of Lemma 4.9 we may assume $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{c},M}\setminus\widetilde{{\mathbf{U}}}_{\beta_{0},M}$. Then by Lemma 4.6 there exists $m_{1}\in\mathbb{N}$ such that $d_{j}\in\left\{\alpha_{(N+1)N},\alpha_{N(N+1)},\alpha_{NN}\right\}$ for all $j>m_{1}$. Now take $m>m_{1}$, and suppose $d_{m+1}\ldots d_{m+3\cdot 2^{n+1}}=\mathbf{a}_{n+1}^{+}$ with $d_{m}^{1}<M$. We will prove
(4.21) $d_{m+3\cdot 2^{n+1}+1}\ldots d_{m+3\cdot 2^{n+2}}=\Phi_{1}(\mathbf{a}_{n+1}^{+})\quad\textrm{or}\quad\Phi_{1}(\mathbf{a}_{n+1}).$
Note by (4.15) and (4.11) that $d_{m+1}^{1}\ldots d_{m+3\cdot 2^{n+1}}^{1}={\mathbf{t}}_{n+1}^{+}={\mathbf{t}}_{n}^{+}\phi_{1}({\mathbf{t}}_{n}).$ Furthermore, $\delta(\beta_{c})$ begins with ${\mathbf{t}}_{n+1}^{+}\phi_{1}({\mathbf{t}}_{n+1})$ by Proposition 4.2. Then by (4.11) and Proposition 2.3 we obtain that
(4.22) $d_{m+3\cdot 2^{n+1}+1}^{1}\ldots d_{m+3\cdot 2^{n+2}}^{1}\preccurlyeq\phi_{1}({\mathbf{t}}_{n+1})=\phi_{1}({\mathbf{t}}_{n}^{+}){\mathbf{t}}_{n}^{+}=\phi_{1}({\mathbf{t}}_{n-1}^{+}){\mathbf{t}}_{n-1}{\mathbf{t}}_{n}^{+}.$
Observe by (4.14) that $d_{m+1}\ldots d_{m+3\cdot 2^{n+1}}=\mathbf{a}_{n+1}^{+}=\mathbf{a}_{n}^{+}\Phi_{1}(\mathbf{a}_{n})=\mathbf{a}_{n}^{+}\Phi_{1}(\mathbf{a}_{n-1}^{+})\mathbf{a}_{n-1}^{+},$ which gives that $d^{1}_{m+3(2^{n}+2^{n-1})}=N<M$ and $d_{m+3(2^{n}+2^{n-1})+1}\ldots d_{m+3\cdot 2^{n+1}}=\mathbf{a}_{n-1}^{+}$. Then by the induction hypothesis for $k=n-1$ we have
$d_{m+3\cdot 2^{n+1}+1}\ldots d_{m+3(2^{n+1}+2^{n-1})}\in\left\{\Phi_{1}(\mathbf{a}_{n-1}^{+}),\Phi_{1}(\mathbf{a}_{n-1})\right\}.$
Note that $\Phi_{1}(\mathbf{a}_{n-1})^{1}=\phi_{1}({\mathbf{t}}_{n-1})\succ\phi_{1}({\mathbf{t}}_{n-1}^{+})$, and $d_{m+3(2^{n}+2^{n-1})}^{1}=N<M$. Then by (4.22) and Proposition 2.3 it follows that
(4.23) $d_{m+3\cdot 2^{n+1}+1}\ldots d_{m+3(2^{n+1}+2^{n-1})}=\Phi_{1}(\mathbf{a}_{n-1}^{+}).$
Note that $d_{m+3\cdot 2^{n+1}}^{1}=N+1>0$. Again by the induction hypothesis we obtain that $d_{m+3(2^{n+1}+2^{n-1})+1}\ldots d_{m+3(2^{n+1}+2^{n})}\in\left\{\mathbf{a}_{n-1},\mathbf{a}_{n-1}^{+}\right\}.$ Then by (4.22) and Proposition 2.3 it follows that
(4.24) $d_{m+3(2^{n+1}+2^{n-1})+1}\ldots d_{m+3(2^{n+1}+2^{n})}=\mathbf{a}_{n-1}.$
So, by (4.23) and (4.24) we have
$d_{m+3\cdot 2^{n+1}+1}\ldots d_{m+3(2^{n+1}+2^{n})}=\Phi_{1}(\mathbf{a}_{n-1}^{+})\mathbf{a}_{n-1}=\Phi_{1}(\mathbf{a}_{n}^{+}).$
Since $d_{m+3\cdot 2^{n+1}}=N+1>0$, by the induction hypothesis for $k=n$ we obtain that $d_{m+3(2^{n+1}+2^{n})+1}\ldots d_{m+3\cdot 2^{n+2}}=\mathbf{a}_{n}$ or $\mathbf{a}_{n}^{+}$. Therefore, by (4.14) we conclude that
$d_{m+3\cdot 2^{n+1}+1}\ldots d_{m+3\cdot 2^{n+2}}=\Phi_{1}(\mathbf{a}_{n}^{+})\mathbf{a}_{n}^{+}=\Phi_{1}(\mathbf{a}_{n+1})\quad\textrm{or}\quad\Phi_{1}(\mathbf{a}_{n}^{+})\mathbf{a}_{n}=\Phi_{1}(\mathbf{a}_{n+1}^{+}),$
establishing (4.21). Hence, by induction this complets the proof. ∎
4.3. Phase transition for $\widetilde{{\mathbf{U}}}_{\beta,M}$ with $M=3N+1$
Recall that the sequence $(\beta_{n})$ is defined in (4.12). Note that $\beta_{n}\nearrow\beta_{c}$ as $n\to\infty$. Our strategy on proving Theorem 1.4 (ii) is to show that the difference set $\widetilde{{\mathbf{U}}}_{\beta_{n+1},M}\setminus\widetilde{{\mathbf{U}}}_{\beta_{n},M}$ is countably for any $n\geq 0$.
Proposition 4.11.
Let $n\in\mathbb{N}_{0}$. Then any sequence in $\widetilde{{\mathbf{U}}}_{\beta_{n+1},M}\setminus\widetilde{{\mathbf{U}}}_{\beta_{n},M}$ must end in
$\bigcup_{j=0}^{2}\left\{\theta^{j}(\mathbf{a}_{n}^{\infty}),\theta^{j}(\Phi_{1}(\mathbf{a}_{n})^{\infty})\right\},$
where $\theta$ is the cyclic permutation defined in (2.2).
First we prove Proposition 4.11 for $n=0$.
Lemma 4.12.
Any sequence in $\widetilde{{\mathbf{U}}}_{\beta_{1},M}\setminus\widetilde{{\mathbf{U}}}_{\beta_{0},M}$ must end in $\mathbf{a}_{0}^{\infty},\Phi_{1}(\mathbf{a}_{0})^{\infty}$ or their cyclic permutations under $\theta$.
Proof.
Take $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{1},M}\setminus\widetilde{{\mathbf{U}}}_{\beta_{0},M}$. By Lemma 4.6 there exists $m\in\mathbb{N}$ such that $d_{m}^{1}<M$ and $d_{m+i}\in\left\{\alpha_{(N+1)N},\alpha_{N(N+1)},\alpha_{NN}\right\}$ for all $i\geq 1.$ Without loss of generality we assume $d_{m+1}=\alpha_{(N+1)N}$. Note that $\delta(\beta_{1})={\mathbf{t}}_{1}^{\infty}=((N+1)N(N+1)NNN)^{\infty}.$ Then by Proposition 2.3 it follows that
(4.25) $d_{m+1}d_{m+2}=\alpha_{(N+1)N}\alpha_{N(N+1)}\quad\textrm{and}\quad d_{m+3}\in\left\{\alpha_{(N+1)N},\alpha_{NN}\right\};$
or
(4.26) $d_{m+1}d_{m+2}=\alpha_{(N+1)N}\alpha_{NN}\quad\textrm{and}\quad d_{m+3}\in\left\{\alpha_{(N+1)N},\alpha_{N(N+1)}\right\}.$
Suppose (4.25) holds, and we assume $d_{m+3}=\alpha_{(N+1)N}$. Then $d_{m+1}d_{m+2}d_{m+3}=\mathbf{a}_{0}^{+}$. By Proposition 4.7 we have $d_{m+4}d_{m+5}d_{m+6}=\Phi_{1}(\mathbf{a}_{0}^{+})$ or $\Phi_{1}(\mathbf{a}_{0})$. However, if $d_{m+4}d_{m+5}d_{m+6}=\Phi_{1}(\mathbf{a}_{0})$, then we have
$d_{m}^{1}<M\quad\textrm{and}\quad d_{m+1}^{1}\ldots d_{m+6}^{1}={\mathbf{t}}_{1}^{+}\succ\delta_{1}(\beta_{1})\ldots\delta_{6}(\beta_{1}),$
leading to a contradiction with $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{1},M}$. So, $d_{m+4}d_{m+5}d_{m+6}=\Phi_{1}(\mathbf{a}_{0}^{+})$. Again, by Proposition 4.7 and the same argument as above we can deduce that
$d_{m+1}d_{m+2}\ldots=(\mathbf{a}_{0}^{+}\Phi_{1}(\mathbf{a}_{0}^{+}))^{\infty}=\mathbf{a}_{1}^{\infty}.$
However, by (4.15) this implies that $d^{1}_{m+1}d^{1}_{m+2}\ldots={\mathbf{t}}_{1}^{\infty}=\delta(\beta_{1})$, which again leads to a contradiction with $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{1},M}$.
In view of (4.25), this implies $d_{m+1}d_{m+2}d_{m+3}=\alpha_{(N+1)N}\alpha_{N(N+1)}\alpha_{NN}=\mathbf{a}_{0}$. By Proposition 2.3 we have that $d_{m+4}=\alpha_{(N+1)N}$ or $\alpha_{N(N+1)}$. If $d_{m+4}=\alpha_{N(N+1)}$, then $d_{m+2}d_{m+3}d_{m+4}=\theta(\mathbf{a}_{0}^{+})$. By the same argument as in the claim we can prove that $d_{m+2}d_{m+3}\ldots=\theta(\mathbf{a}_{1}^{\infty})$, which will lead to a contradiction. So, $d_{m+4}=\alpha_{(N+1)N}$. Then by Proposition 2.3 it follows that $d_{m+5}=\alpha_{NN}$ or $\alpha_{N(N+1)}$. If $d_{m+5}=\alpha_{NN}$, then $d_{m+3}d_{m+4}d_{m+5}=\theta^{2}(\mathbf{a}_{0}^{+})$, again the same argument as in the claim yields that $d_{m+3}d_{m+4}\ldots=\theta^{2}(\mathbf{a}_{1}^{\infty})$, which will lead to a contradiction. Therefore, $d_{m+5}=\alpha_{N(N+1)}$. Repeating the above argument we conclude that $d_{m+1}d_{m+2}\ldots=\mathbf{a}_{0}^{\infty}$.
Similarly, if (4.26) holds, then by the same argument as above we could prove that
$d_{m+1}d_{m+2}\ldots=\theta(\Phi_{1}(\mathbf{a}_{0}))^{\infty},$
completing the proof. ∎
Proof of Proposition 4.11.
We will prove the theorem by induction on $n$. Note by Lemma 4.12 that it holds for $n=0$. Now suppose the result holds for all $n\leq k$ for some $k\geq 0$, and we consider it for $n=k+1$. Let $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{k+2},M}\setminus\widetilde{{\mathbf{U}}}_{\beta_{k+1},M}$. Then by Lemma 4.6 there exists $m_{1}\in\mathbb{N}$ such that $d_{m_{1}}^{1}<M$ and $d_{m_{1}+i}\in\left\{\alpha_{(N+1)N},\alpha_{N(N+1)},\alpha_{NN}\right\}$ for all $i\geq 1.$ Note by the induction hypothesis that any sequence in $\widetilde{{\mathbf{U}}}_{\beta_{k+1},M}\setminus\widetilde{{\mathbf{U}}}_{\beta_{0},M}=\bigcup_{j=0}^{k}(\widetilde{{\mathbf{U}}}_{\beta_{j+1},M}\setminus\widetilde{{\mathbf{U}}}_{\beta_{j},M})$ must end with $\mathbf{a}_{j}^{\infty},\Phi_{1}(\mathbf{a}_{j}^{\infty})$ or their cyclic permutations under $\theta$ for some $0\leq j\leq k$. Observe by (4.15) that $\mathbf{a}_{j}^{1}={\mathbf{t}}_{j}$ for all $j\geq 0$, and ${\mathbf{t}}_{0}^{\infty}\prec{\mathbf{t}}_{1}^{\infty}\prec\cdots\prec{\mathbf{t}}_{k}^{\infty}\prec{\mathbf{t}}_{k+1}^{\infty}$. Since $\delta(\beta_{j})={\mathbf{t}}_{j}^{\infty}$ for all $j\geq 0$, by (4.14) it follows that the sequence $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{k+2},M}\setminus\widetilde{{\mathbf{U}}}_{\beta_{k+1},M}$ must contain a block in $\bigcup_{j=0}^{2}\left\{\theta^{j}(\mathbf{a}_{k}^{+}),\theta^{j}(\Phi_{1}(\mathbf{a}_{k}^{+}))\right\}$. Without loss of generality we may assume $d_{m_{2}+1}\ldots d_{m_{2}+3\cdot 2^{k}}=\mathbf{a}_{k}^{+}$ for some $m_{2}\geq m_{1}$. Note that $d^{1}_{m_{2}}<M$. Then by Proposition 4.7 it follows that
(4.27) $d_{m_{2}+1}\ldots d_{m_{2}+3\cdot 2^{k+1}}=\mathbf{a}_{k}^{+}\Phi_{1}(\mathbf{a}_{k}^{+})=\mathbf{a}_{k+1}\quad\textrm{or}\quad\mathbf{a}_{k}^{+}\Phi(\mathbf{a}_{k})=\mathbf{a}_{k+1}^{+}.$
If $d_{m_{2}+1}\ldots d_{m_{2}+3\cdot 2^{k+1}}=\mathbf{a}_{k+1}^{+}$, the by Proposition 4.7 and using $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{k+2},M}$ it follows that $d_{m_{2}+1}d_{m_{2}+2}\ldots=\mathbf{a}_{k+2}^{\infty}$. But by (4.15) this gives
$d_{m_{2}+1}^{1}d_{m_{2}+2}^{1}\ldots={\mathbf{t}}_{k+2}^{\infty}=\delta(\beta_{k+2}),$
which leads to a contradiction with $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{k+2},M}$.
Hence, by (4.27) we must have $d_{m_{2}+1}\ldots d_{m_{2}+3\cdot 2^{k+1}}=\mathbf{a}_{k+1}=\mathbf{a}_{k}^{+}\Phi_{1}(\mathbf{a}_{k}^{+})$. Again, by Proposition 4.7 it follows that
(4.28) $d_{m_{2}+3\cdot 2^{k+1}+1}\ldots d_{m_{2}+3(2^{k+1}+2^{k})}=\mathbf{a}_{k}\quad\textrm{or}\quad\mathbf{a}_{k}^{+}.$
If $d_{m_{2}+3\cdot 2^{k+1}+1}\ldots d_{m_{2}+3(2^{k+1}+2^{k})}=\mathbf{a}_{k}$, then $d_{m_{2}+3\cdot 2^{k}+1}\ldots d_{m_{2}+3\cdot(2^{k+1}+2^{k})}=\Phi_{1}(\mathbf{a}_{k}^{+})\mathbf{a}_{k}=\Phi(\mathbf{a}_{k+1}^{+})$. By (4.15), Proposition 2.3 and using $d^{1}_{m_{2}+3\cdot 2^{k}}>0$ it follows that
$\overline{d^{\oplus}_{m_{2}+3\cdot 2^{k}+1}}\,\overline{d^{\oplus}_{m_{2}+3\cdot 2^{k}+2}}\ldots\succcurlyeq{\mathbf{t}}_{k+2}^{\infty}=\delta(\beta_{k+2}),$
again leading to a contradiction with $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{k+2},M}$.
So, by (4.28) we have $d_{m_{2}+3\cdot 2^{k+1}+1}\ldots d_{m_{2}+3(2^{k+1}+2^{k})}=\mathbf{a}_{k}^{+}$. Continuing this argument as in (4.27) and (4.28) we can prove that $d_{m_{2}+1}d_{m_{2}+2}\ldots=(\mathbf{a}_{k}^{+}\Phi_{1}(\mathbf{a}_{k}^{+}))^{\infty}=\mathbf{a}_{k+1}^{\infty}.$ This proves the result for $n=k+1$. ∎
Proof of Theorem 1.4 for $M=3N+1$.
By Propositions 2.5, 3.1 and 4.2 it suffices to prove (ii)–(iv) for $\widetilde{U}_{\beta,M}$. Since the map $\Pi_{\beta}$ from $\widetilde{{\mathbf{U}}}_{\beta,M}$ to $\widetilde{U}_{\beta,M}$ is bijective and bi-Hölder continuous, we only need to consider the symbolic intrinsic univoque set $\widetilde{{\mathbf{U}}}_{\beta,M}$. Here we equip with the symbolic space ${\Omega}^{\mathbb{N}}_{M}$ the metric $\rho$ given by
(4.29) $\rho((c_{i}),(d_{i}))=2^{-\inf\left\{n:c_{n}\neq d_{n}\right\}}$
for any $(c_{i}),(d_{i})\in{\Omega}_{M}^{\mathbb{N}}$.
First we prove (ii). Take $\beta\in(\beta_{G},\beta_{c})$. Note that $\beta_{0}=\beta_{G}$ and $\beta_{k}\nearrow\beta_{c}$ as $k\to\infty$. So there exists $k\in\mathbb{N}$ such that $\beta<\beta_{k}$. Note by Propositions 3.1 and 4.11 that $\widetilde{{\mathbf{U}}}_{\beta_{k},M}=\widetilde{{\mathbf{U}}}_{\beta_{0},M}\cup\bigcup_{j=0}^{k-1}\widetilde{{\mathbf{U}}}_{\beta_{j+1},M}\setminus\widetilde{{\mathbf{U}}}_{\beta_{j},M}$ is at most countable. This implies that $\widetilde{{\mathbf{U}}}_{\beta,M}\subset\widetilde{{\mathbf{U}}}_{\beta_{k},M}$ is at most countable. On the other hand, by Proposition 2.3 we have
$\alpha_{00}^{n}(\alpha_{(N+1)N}\alpha_{N(N+1)}\alpha_{NN})^{\infty}\in\widetilde{{\mathbf{U}}}_{\beta,M}\quad\forall n\in\mathbb{N}.$
So, $\widetilde{{\mathbf{U}}}_{\beta,M}$ is indeed countably infinite, establishing (ii).
Next we prove (iii). The proof is similar to [23, Proposition 5.6]. Note that $\beta_{k}\nearrow\beta_{c}$ as $k\to\infty$. By Proposition 4.7 it follows that $\widetilde{{\mathbf{U}}}_{\beta_{c},M}$ contains all of the following sequences
$\mathbf{a}_{1}^{j_{1}}\mathbf{a}_{2}^{j_{2}}\cdots\mathbf{a}_{k}^{j_{k}}\cdots,\quad j_{k}\in\mathbb{N}.$
Note that each block $\mathbf{a}_{k}$ can not be written as concatenations of two or more blocks of the form $\mathbf{a}_{l}$ with $l<k$. This implies that $\widetilde{{\mathbf{U}}}_{\beta_{c},M}$ is uncountable. On the other hand, by the recursive construction as in the proof of Proposition 4.7 it follows that any sequence in $\widetilde{{\mathbf{U}}}_{\beta_{c},M}\setminus\widetilde{{\mathbf{U}}}_{\beta_{G},M}$ must be of the form
$\omega(\mathbf{a}_{i_{1}}^{+}\Phi_{1}(\mathbf{a}_{i_{1}}^{+}))^{j_{1}}\left(\mathbf{a}_{i_{1}}^{+}\Phi_{1}(\mathbf{a}_{i_{2}^{\prime}}^{+})\right)^{j_{1}^{\prime}}\cdots(\mathbf{a}_{i_{k}}^{+}\Phi_{1}(\mathbf{a}_{i_{k}}^{+}))^{j_{k}}\left(\mathbf{a}_{i_{k}}^{+}\Phi_{1}(\mathbf{a}_{i_{k+1}^{\prime}}^{+})\right)^{j_{k}^{\prime}}\cdots$
or its cyclic permutations under $\theta$, where $\omega\in\Omega_{M}^{*}$, $j_{k}\in\mathbb{N}\cup\left\{0,\infty\right\}$, $j_{k}^{\prime}\in\left\{0,1\right\}$ and
$1\leq i_{1}<i_{2}^{\prime}\leq i_{2}<\cdots<i_{k}^{\prime}\leq i_{k}<i_{k+1}^{\prime}\leq\cdots.$
Observe that the length of $\mathbf{a}_{k}$ grows exponentially fast. This implies that $\dim_{H}\widetilde{{\mathbf{U}}}_{\beta_{c},M}=0$, establishing (iii).
Finally we prove (iv). For $k\in\mathbb{N}_{0}$ note that ${\mathbf{t}}_{k}^{+}=\lambda_{1}\ldots\lambda_{3\cdot 2^{k}}$. Then by Lemma 4.3 it follows that $\sigma^{i}({\mathbf{t}}_{k}^{+}\phi_{1}({\mathbf{t}}_{k})^{\infty})\preccurlyeq{\mathbf{t}}_{k}^{+}\phi_{1}({\mathbf{t}}_{k})^{\infty}$ for all $i\geq 0$. Then by Lemma 2.1 there exists $\hat{\beta}_{k}\in(N+1,N+2)$ such that
$\delta(\hat{\beta}_{k})={\mathbf{t}}_{k}^{+}\phi_{1}({\mathbf{t}}_{k})^{\infty}.$
Since $\delta(\hat{\beta}_{k})\searrow\delta(\beta_{c})$ as $k\to\infty$, again by Lemma 2.1 it follows that $\hat{\beta}_{k}\searrow\beta_{c}$ as $k\to\infty$. So, for $\beta>\beta_{c}$ there exists a smallest $k\in\mathbb{N}_{0}$ such that $\hat{\beta}_{k}<\beta$. Then we can show that $\widetilde{{\mathbf{U}}}_{\beta,M}$ contains a subshift of finite type $(X_{k},\sigma)$ represented by the labeled graph $G$ (see Figure 5).
$\Phi_{1}(\mathbf{a}_{k})$$\Phi_{1}(\mathbf{a}_{k}^{+})$$\mathbf{a}_{k}$$\mathbf{a}_{k}^{+}$
Figure 5. The labeled graph $G$ presenting $(X_{k},\sigma)$.
From this we can deduce that
$\dim_{H}\widetilde{{\mathbf{U}}}_{\beta,M}\geq\dim_{H}X_{k}=\frac{\log 2}{|\mathbf{a}_{k}|\log 2}=\frac{1}{|\mathbf{a}_{k}|}>0,$
where we use the metric $\rho$ defined in (4.29). This completes the proof. ∎
At the end of this section we give an example to illustrate that the Hausdorff dimension of $\widetilde{U}_{\beta,M}$ can be explicitly calculated for some $\beta>\beta_{c}(M)$.
Example 4.13.
Let $\beta=\frac{1+\sqrt{5}}{2}$ and $M=1$. Then $\beta>\beta_{c}(1)$. Observe by Proposition 2.3 that up to countably many points the set $\widetilde{{\mathbf{U}}}_{\beta,1}$ is a subshift of finite type in $\left\{\alpha_{00},\alpha_{10},\alpha_{01}\right\}^{\mathbb{N}}$ with the set of forbidden words
=$\mathscr{F}=\left\{\alpha_{10}\alpha_{10},\alpha_{01}\alpha_{01},\alpha_{00}\alpha_{00}\right\}.$
Then it can be represented as a directed graph with the vertex set $V=\left\{\alpha_{00},\alpha_{10},\alpha_{01}\right\}$ and the adjacency matrix
$A=\left(\begin{array}[]{ccc}0&1&1\\ 1&0&1\\ 1&1&0\end{array}\right).$
So, $\dim_{H}\widetilde{U}_{\beta,1}=\frac{\log\rho(A)}{\log\beta}=\frac{\log 2}{\log\beta}\approx 1.44042$, where $\rho(A)$ denotes the spectral radius of $A$. Indeed, we also have $\dim_{H}U_{\beta,1}=\frac{\log 2}{\log\beta}$.
We remark that if $\beta$ is a mutinacci number, then by a similar argument as in Example 4.13 we are able to determine the explicit Hausdorff dimension of $U_{\beta,1}$.
5. Generalized Komornik-Loreti constant $\beta_{c}(M)$ for $M=3N+2$
Let $M=3N+2$ with $N\in\mathbb{N}_{0}$. Recall $\beta_{c}(M)$ from Definition 1.2. We will show in this section that $\beta_{c}(M)$ is the critical base for $\widetilde{{\mathbf{U}}}_{\beta,M}$, i.e., $\widetilde{{\mathbf{U}}}_{\beta,M}$ has positive Hausdorff dimension if and only if $\beta>\beta_{c}(M)$. The proof is similar to that for $M=3N+1$, so we only sketch the main idea in the proof of Theorem 1.4 for $M=3N+2$.
First we introduce a Thue-Morse type sequence $(\gamma_{i})\in\left\{N,N+1\right\}^{\mathbb{N}}$, which will be the quasi-greedy $\beta_{c}(M)$-expansion of $1$. In view of Definition 1.2, we define the sequence $(\gamma_{i})$ by
(5.1) $\gamma_{3k+1}=\tau_{2k+1}+N,\quad\gamma_{3k+2}=N+1,\quad\gamma_{3k+3}=\tau_{2k+2}+N,$
[…]
$(\gamma_{i})_{i=1}^{\infty}=\sigma(\varphi_{2}(\tau_{0})\varphi_{2}(\tau_{1})\varphi_{2}(\tau_{2})\ldots),$
where $\sigma$ is the left-shift map.
Similar to $(\lambda_{i})$ defined in the previous section, we construct a sequence of blocks $({\mathbf{u}}_{n})$ approximating $(\gamma_{i})$ componentwisely. Let
${\Sigma}_{2}:=\left\{(N+1)(N+1)(N+1),N(N+1)N,(N+1)(N+1)N,N(N+1)(N+1)\right\},$
and we define a block map $\phi_{2}:{\Sigma}_{2}\rightarrow{\Sigma}_{2}$ by
$\begin{split}(N+1)(N+1)(N+1)\quad&\longleftrightarrow\quad N(N+1)N,\\ (N+1)(N+1)N\quad&\longleftrightarrow\quad N(N+1)(N+1).\end{split}$
Then the sequence $({\mathbf{u}}_{n})$ is defined recursively as follows:
(5.2) ${\mathbf{u}}_{0}=(N+1)(N+1)N,\quad\textrm{and}\quad{\mathbf{u}}_{n+1}:={\mathbf{u}}_{n}^{+}\phi_{2}({\mathbf{u}}_{n}^{+})\quad\forall n\geq 0.$
For example,
${\mathbf{u}}_{1}$ $=(N+1)(N+1)(N+1)N(N+1)N,$
${\mathbf{u}}_{2}$ $=(N+1)(N+1)(N+1)N(N+1)(N+1)\,N(N+1)N(N+1)(N+1)N.$
Note that $(\gamma_{i})$ begins with ${\mathbf{u}}_{n}^{+}$ for any $n\geq 1$. Then the sequence $({\mathbf{u}}_{n})$ converges to $(\gamma_{i})$ componentwisely, i.e.,
$(\gamma_{i})=\lim_{n\to\infty}{\mathbf{u}}_{n}\in\Sigma_{2}^{\mathbb{N}}.$
$x$$y$$\alpha_{00}$$\alpha_{10}$$\alpha_{20}$$\alpha_{01}$$\alpha_{02}$$\alpha_{11}$$\frac{1}{\beta}$$\frac{2}{\beta}$$\frac{2}{\beta-1}$$\frac{1}{\beta}$$\frac{2}{\beta}$$\frac{2}{\beta-1}$
Figure 6. The graph for the first generation of $S_{\beta,M}$ with $M=2$ and $\beta=2.5$. The convex hull $\Delta_{\beta,M}$ is the triangle with vertices $(0,0),(2/(\beta-1),0)$ and $(0,2/(\beta-1))$. Each light grey triangle corresponds to a $f_{\beta,\alpha_{ij}}(\Delta_{\beta,M})$ for some $\alpha_{ij}\in\Omega_{M}$. Furthermore, the overlap region $O_{\beta,M}$ is the union of seven small dark grey triangles.
We also define the conjugacy of $(\gamma_{i})$ by
$(\hat{\gamma}_{i})_{i=1}^{\infty}:=\phi_{2}(\gamma_{1}\gamma_{2}\gamma_{3})\phi_{2}(\gamma_{4}\gamma_{5}\gamma_{6})\ldots\phi_{2}(\gamma_{3k+1}\gamma_{3k+2}\gamma_{3k+3})\ldots.$
Then by (5.1) and (5.2) it follows that for any $k\in\mathbb{N}_{0}$,
(5.3) $\hat{\gamma}_{3k+1}=\overline{\tau_{2k+1}}+N,\quad\hat{\gamma}_{3k+2}=N+1,\quad\hat{\gamma}_{3k+3}=\overline{\tau_{2k+2}}+N,$
where $\overline{\tau_{i}}=1-\tau_{i}$ for any $i\geq 0$.
By a similar argument as in the proof of Lemma 4.3 we have the following result.
Lemma 5.1.
For any $n\in\mathbb{N}_{0}$ and any $0\leq i<3\cdot 2^{n}$, the following inequalities hold:
$\hat{\gamma}_{1}\ldots\hat{\gamma}_{3\cdot 2^{n}-i}$ $\prec\gamma_{i+1}\ldots\gamma_{3\cdot 2^{n}}\preccurlyeq\gamma_{1}\ldots\gamma_{3\cdot 2^{n}-i},$
$\hat{\gamma}_{1}\ldots\hat{\gamma}_{3\cdot 2^{n}-i}$ $\preccurlyeq\hat{\gamma}_{i+1}\ldots\hat{\gamma}_{3\cdot 2^{n}}\prec\gamma_{1}\ldots\gamma_{3\cdot 2^{n}-i}.$
By Lemmas 5.1 and 2.1 it follows that $(\gamma_{i})$ is the quasi-greedy $\beta_{c}(M)$-expansion of $1$, i.e.,
$\begin{split}\delta(\beta_{c}(M))&=\gamma_{1}\gamma_{2}\ldots=(N+1)(N+1)(N+1)N(N+1)(N+1)\;N(N+1)N\ldots,\end{split}$
By Lemma 4.4, (5.1) and the same argument as in the proof of Proposition 4.2 we can prove that $\beta_{c}(3N+2)$ is transcendental.
Note by (5.2) that ${\mathbf{u}}_{n}=\gamma_{1}\ldots\gamma_{3\cdot 2^{n}}^{-}$. Then by Lemma 5.1 we have $\sigma^{i}({\mathbf{u}}_{n}^{\infty})\preccurlyeq{\mathbf{u}}_{n}^{\infty}$ for all $i\geq 0.$ So, by Lemma 2.1 there exists a unique $\beta_{n}=\beta_{n}(M)\in(1,M+1]$ such that
(5.4) $\delta(\beta_{n})={\mathbf{u}}_{n}^{\infty}.$
Note by Proposition 3.1 that $\delta(\beta_{G}(M))=((N+1)(N+1)N)^{\infty}={\mathbf{u}}_{0}^{\infty}=\delta(\beta_{0}(M)).$ Then $\beta_{G}(M)=\beta_{0}(M)$ by Lemma 2.1. Observe that $\delta(\beta_{n})={\mathbf{u}}_{n}^{\infty}\nearrow(\gamma_{i})=\delta(\beta_{c})$ as $n\to\infty$. By Lemma 2.1 it follows that
$\beta_{G}=\beta_{0}<\beta_{1}<\cdots<\beta_{n}<\beta_{n+1}<\cdots,\quad\textrm{and}\quad\beta_{n}\nearrow\beta_{c}\quad\textrm{as }n\to\infty.$
To prove Theorem 1.4 for $M=3N+2$ we will construct a sequence of blocks $\left\{{\mathbf{b}}_{n}\right\}_{n=0}^{\infty}$ in ${\Omega}_{M}^{*}$. Let
$\pazocal A_{2}$ :=$:=\left\{\alpha_{(N+1)(N+1)}\alpha_{(N+1)N}\alpha_{N(N+1)},\quad\alpha_{N(N+1)}\alpha_{(N+1)N}\alpha_{(N+1)(N+1)},\right.$
$\hskip 21.33955pt\left.\alpha_{(N+1)(N+1)}\alpha_{(N+1)N}\alpha_{(N+1)(N+1)},\quad\alpha_{N(N+1)}\alpha_{(N+1)N}\alpha_{N(N+1)}\right\}.$
We define the substitution $\Phi_{2}:\pazocal A_{2}\to\pazocal A_{2}$ by
(5.5) $\begin{split}\alpha_{(N+1)(N+1)}\alpha_{(N+1)N}\alpha_{N(N+1)}\quad&\longleftrightarrow\quad\alpha_{N(N+1)}\alpha_{(N+1)N}\alpha_{(N+1)(N+1)},\\ \alpha_{(N+1)(N+1)}\alpha_{(N+1)N}\alpha_{(N+1)(N+1)}\quad&\longleftrightarrow\quad\alpha_{N(N+1)}\alpha_{(N+1)N}\alpha_{N(N+1)}.\end{split}$
Then the sequence $\left\{{\mathbf{b}}_{n}\right\}_{n=0}^{\infty}\subset\pazocal A_{1}^{*}$ is defined as follows (see Figure 6):
(5.6) ${\mathbf{b}}_{0}=\alpha_{(N+1)(N+1)}\alpha_{(N+1)N}\alpha_{N(N+1)},\quad\textrm{and}\quad{\mathbf{b}}_{n+1}={\mathbf{b}}_{n}^{+}\Phi_{2}({\mathbf{b}}_{n}^{+})\quad\forall n\geq 0,$
where ${\mathbf{b}}_{n}^{+}$ is the word in $\pazocal A_{2}^{*}$ by changing the last digit of ${\mathbf{b}}_{n}$ from $\alpha_{N(N+1)}$ to $\alpha_{(N+1)(N+1)}$. Then each ${\mathbf{b}}_{n}$ has length $3\cdot 2^{n}$ and ends with $\alpha_{N(N+1)}$. If we write ${\mathbf{b}}_{n}=b_{1}\ldots b_{3\cdot 2^{n}}$, then $b_{3k+2}=\alpha_{(N+1)N}$ for all $0\leq k<2^{n}$. So,
${\mathbf{b}}_{0}\sim\left(\begin{array}[]{ccc}N+1&N+1&N\\ N+1&N&N+1\\ N&N+1&N+1\end{array}\right)$
and
${\mathbf{b}}_{1}\sim\left(\begin{array}[]{cccccc}N+1&N+1&N+1&N&N+1&N\\ N+1&N&N+1&N+1&N&N+1\\ N&N+1&N&N+1&N+1&N+1\end{array}\right).$
Comparing the definitions of $({\mathbf{u}}_{n})$ and $({\mathbf{b}}_{n})$ in (5.2) and (5.6) respectively, it follows that
(5.7) ${\mathbf{b}}_{n}^{1}={\mathbf{u}}_{n},\quad{\mathbf{b}}_{n}^{2}=((N+1)N(N+1))^{n}\quad\textrm{and}\quad\overline{{\mathbf{b}}_{n}^{\oplus}}=\phi_{2}({\mathbf{u}}_{n})$
for any $n\geq 0$.
The following result can be proved in a similar way as that for Proposition 4.11.
Proposition 5.2.
Let $n\in\mathbb{N}_{0}$. Then any sequence in $\widetilde{{\mathbf{U}}}_{\beta_{n+1},M}\setminus\widetilde{{\mathbf{U}}}_{\beta_{n},M}$ must end in
$\bigcup_{j=0}^{2}\left\{\theta^{j}({\mathbf{b}}_{n}^{\infty}),\theta^{j}(\Phi_{2}({\mathbf{b}}_{n})^{\infty})\right\},$
where $\theta$ is the cyclic permutation defined in (2.2).
It is worth mentioning that for $M=3N+2$, if $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{c},M}\setminus\widetilde{{\mathbf{U}}}_{\beta_{0},M}$, then there exists $m\in\mathbb{N}$ such that $d_{m}^{1}<M,d_{m}^{2}<M,\overline{d_{m}^{\oplus}}<M$, and
$d_{k}\in\bigcup_{j=0}^{2}\left\{\theta^{j}(\alpha_{(N+1)(N+1)})\right\}\quad\forall k>m.$
So, in this case we don’t need a lemma like Lemma 4.5. All other properties are similar to that for $M=3N+1$. For example, we have a similar characterization on patterns appearing in sequences of $\widetilde{{\mathbf{U}}}_{\beta_{c},M}$.
Proposition 5.3.
Let $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{c},M}$. Then the following statements hold.
  1. (i)
    If $d^{1}_{m}<M$ and $d_{m+1}\ldots d_{m+3\cdot 2^{n}}={\mathbf{b}}_{n}^{+}$, then $d_{m+3\cdot 2^{n}+1}\ldots d_{m+3\cdot 2^{n+1}}\in\left\{\Phi_{1}({\mathbf{b}}_{n}),\Phi_{1}({\mathbf{b}}_{n}^{+})\right\}.$
  2. (ii)
    If $d^{1}_{m}>0$ and $d_{m+1}\ldots d_{m+3\cdot 2^{n}}=\Phi_{1}({\mathbf{b}}_{n}^{+})$, then $d_{m+3\cdot 2^{n}+1}\ldots d_{m+3\cdot 2^{n+1}}\in\left\{{\mathbf{b}}_{n},{\mathbf{b}}_{n}^{+}\right\}.$
The proof of Proposition 5.3 is very similar to Proposition 4.7, and we leave it as an exercise for interested readers.
Proof of Theorem 1.4 for $M=3N+2$.
The proof is similar to that for $M=3N+1$. By using Proposition 5.2 we can deduce that for $\beta<\beta_{c}$ the set $\widetilde{{\mathbf{U}}}_{\beta,M}$ is at most countable; and by Proposition 5.3 we can prove that for $\beta=\beta_{c}$ the set $\widetilde{{\mathbf{U}}}_{\beta,M}$ is uncountable but has zero Hausdorff dimension. Furthermore, for $\beta>\beta_{c}$ we can construct a subshift of finite type $Y_{k}\subset\widetilde{{\mathbf{U}}}_{\beta,M}$ having positive Hausdorff dimension, where $Y_{k}$ is analogous to $X_{k}$ as described in Figure 5 with $\mathbf{a}_{k},\mathbf{a}_{k}^{+},\Phi_{1}(\mathbf{a}_{k})$ and $\Phi_{1}(\mathbf{a}_{k}^{+})$ replaced by ${\mathbf{b}}_{k},{\mathbf{b}}_{k}^{+},\Phi_{2}({\mathbf{b}}_{k})$ and $\Phi_{2}({\mathbf{b}}_{k}^{+})$, respectively. ∎
6. Generalized Komornik-Loreti constant for $\widetilde{{\mathbf{U}}}_{\beta,M}$ with $M=3N+3$
Let $M=3N+3$ with $N\in\mathbb{N}_{0}$. Recall that $\beta_{c}(M)$ is defined in Definition 1.2. In this section we will prove Theorem 1.4 for $M=3N+3$ that $\widetilde{{\mathbf{U}}}_{\beta,M}$ has positive Hausdorff dimension if and only if $\beta>\beta_{c}(M)$. Completely different from the case for $M=1$, we need more effort to find the critical base $\beta_{c}(M)$ for $M=3N+3$.
In view of Definition 1.2 and the definition of the classical Thue-Morse sequence $(\tau_{i})$, we introduce a Thue-Morse type sequence $(\eta_{i})\in\left\{N,N+1,N+2\right\}^{\mathbb{N}}$:
(6.1) $\eta_{3k+1}=N+2-2\tau_{k},\quad\eta_{3k+2}=N+\tau_{k},\quad\eta_{3k+3}=N+1+\tau_{k+1}\quad\forall k\in\mathbb{N}_{0}.$
[…]
$(\eta_{i})_{i=1}^{\infty}=\sigma(\varphi_{3}(\tau_{0})\varphi_{3}(\tau_{1})\varphi_{3}(\tau_{2})\ldots),$
where $\sigma$ is the left shift map.
Proposition 6.1.
Let $M=3N+3$ with $N\in\mathbb{N}_{0}$. Then the quasi-greedy $\beta_{c}(M)$-expansion of $1$ is given by
$\delta(\beta_{c}(M))=\eta_{1}\eta_{2}\ldots=(N+2)N(N+2)N(N+1)(N+2)N(N+1)(N+1)(N+2)N(N+1)\ldots.$
Furthermore, $\beta_{c}(M)$ is transcendental.
Let
${\Sigma}_{3}:=\left\{(N+2)N(N+2),N(N+1)(N+1),(N+2)N(N+1),N(N+1)(N+2)\right\},$
and we define the block map $\phi_{3}:{\Sigma}_{3}\rightarrow{\Sigma}_{3}$ by
(6.2) $\begin{split}(N+2)N(N+2)\quad&\longleftrightarrow\quad N(N+1)(N+1),\\ (N+2)N(N+1)\quad&\longleftrightarrow\quad N(N+1)(N+2).\end{split}$
Accordingly, we define the conjugacy of $(\eta_{i})$ by
$(\hat{\eta}_{i})_{i=1}^{\infty}:=\phi_{3}(\eta_{1}\eta_{2}\eta_{3})\phi_{3}(\eta_{4}\eta_{5}\eta_{6})\ldots\phi_{3}(\eta_{3k+1}\eta_{3k+2}\eta_{3k+3})\ldots.$
Then $(\hat{\eta}_{i})\in\Sigma_{3}^{\mathbb{N}}$ begins with $N(N+1)(N+1)(N+2)N(N+1)$. Furthermore, by the definition of $\phi_{3}$ it follows that for any $k\in\mathbb{N}_{0}$,
(6.3) $\hat{\eta}_{3k+1}=N+2-2\overline{\tau_{k}},\quad\hat{\eta}_{3k+2}=N+\overline{\tau_{k}},\quad\hat{\eta}_{3k+3}=N+1+\overline{\tau_{k+1}},$
where $\overline{\tau_{i}}=1-\tau_{i}$ for all $i\geq 0$.
Lemma 6.2.
For any $n\in\mathbb{N}_{0}$ and any $0\leq i<3\cdot 2^{n}$, the following inequalities hold:
(6.4) $\hat{\eta}_{1}\ldots\hat{\eta}_{3\cdot 2^{n}-i}$ $\prec\eta_{i+1}\ldots\eta_{3\cdot 2^{n}}\preccurlyeq\eta_{1}\ldots\eta_{3\cdot 2^{n}-i},$
(6.5) $\hat{\eta}_{1}\ldots\hat{\eta}_{3\cdot 2^{n}-i}$ $\preccurlyeq\hat{\eta}_{i+1}\ldots\hat{\eta}_{3\cdot 2^{n}}\prec\eta_{1}\ldots\eta_{3\cdot 2^{n}-i}.$
Proof.
Since the proof of (6.5) is similar, we only prove (6.4). Take $i\in\left\{0,1,\ldots,3\cdot 2^{n}-1\right\}$. We consider the following three cases: (I) $i=3k$, (II) $i=3k+1$, and (III) $i=3k+2$ for some $k\in\left\{0,1,\ldots,2^{n}-1\right\}$.
Case (I). $i=3k$ with $0\leq k<2^{n}$. If $k=2^{n}-1$, then by (6.1) we have
$\eta_{i+1}\eta_{i+2}\eta_{i+3}$ $=(N+2-2\tau_{2^{n}-1})(N+\tau_{2^{n}-1})(N+1+\tau_{2^{n}})$
$\preccurlyeq(N+2)N(N+2)=\eta_{1}\eta_{2}\eta_{3}.$
On the other hand, by (6.3) and using $\tau_{2^{n}}=1$ it follows that
$\eta_{i+1}\eta_{i+2}\eta_{i+3}\succcurlyeq N(N+1)(N+2)\succ N(N+1)(N+1)=\hat{\eta}_{1}\hat{\eta}_{2}\hat{\eta}_{3}.$
So, in the following we assume $10\leq k<2^{n}-$ 1. Note by (6.1) that
(6.6) $\eta_{i+1}\ldots\eta_{3\cdot 2^{n}}=(N+2-2\tau_{k})(N+\tau_{k})\mathbf{B}(\tau_{k+1})\cdots\mathbf{B}(\tau_{2^{n}-1})(N+1+\tau_{2^{n}}),$
where
$\mathbf{B}(\tau_{j}):=(N+1+\tau_{j})(N+2-2\tau_{j})(N+\tau_{j})\quad\textrm{for }j\geq 1.$
Observe by (6.1) and (6.3) that
(6.7) $\begin{split}\eta_{1}\ldots\eta_{3\cdot 2^{n}}&=(N+2)N\mathbf{B}(\tau_{1})\cdots\mathbf{B}(\tau_{2^{n}-1})(N+1+\tau_{2^{n}}),\\ \hat{\eta}_{1}\ldots\hat{\eta}_{3\cdot 2^{n}}&=N(N+1)\mathbf{B}(\overline{\tau_{1}})\cdots\mathbf{B}(\overline{\tau_{2^{n}-1}})(N+1+\overline{\tau_{2^{n}}}).\end{split}$
Since $\mathbf{B}(\tau_{i})\prec\mathbf{B}(\tau_{j})$ is equivalent to $\tau_{i}<\tau_{j}$, by (6.6) and (6.7) it follows that
$\quad\eta_{i+1}\ldots\eta_{3\cdot 2^{n}}(N+2-2\tau_{2^{n}})(N+\tau_{2^{n}})$
$\preccurlyeq(N+2)N\mathbf{B}(\tau_{k+1})\cdots\mathbf{B}(\tau_{2^{n}})$
$\preccurlyeq(N+2)N\mathbf{B}(\tau_{1})\cdots\mathbf{B}(\tau_{2^{n}-k})=\eta_{1}\ldots\eta_{3\cdot 2^{n}-i+2},$
where the second inequality follows by using the property of $(\tau_{i})$ that $\tau_{k+1}\ldots\tau_{2^{n}}\preccurlyeq\tau_{1}\ldots\tau_{2^{n}-k}$. This proves the second inequality in (6.4).
On the other hand, by (6.6) and (6.7) we obtain that
(6.8) $\begin{split}&\quad\eta_{i+1}\ldots\eta_{3\cdot 2^{n}}(N+2-2\tau_{2^{n}})(N+\tau_{2^{n}})\\ &\succcurlyeq N(N+1)\mathbf{B}(\tau_{k+1})\cdots\mathbf{B}(\tau_{2^{n}})\\ &\succ N(N+1)\mathbf{B}(\overline{\tau_{1}})\cdots\mathbf{B}(\overline{\tau_{2^{n}-k}})=\hat{\eta}_{1}\ldots\hat{\eta}_{3\cdot 2^{n}-i+2},\end{split}$
where the second inequality follows by using the fact that $\tau_{k+1}\ldots\tau_{2^{n}}\succ\overline{\tau_{1}}\ldots\overline{\tau_{2^{n}-k}}$. Note that $(N+2-2\tau_{2^{n}})(N+\tau_{2^{n}})=N(N+1)\preccurlyeq\hat{\eta}_{j}\hat{\eta}_{j+1}$ for all $j\geq 1$. So, by (6.8) it follows that
$\eta_{i+1}\ldots\eta_{3\cdot 2^{n}}\succ\hat{\eta}_{1}\ldots\hat{\eta}_{3\cdot 2^{n}-i},$
establishing (6.4) when $i=3k$.
Case (II). $i=3k+1$ with $0\leq k<2^{n}$. Then $\eta_{i+1}=N+\tau_{k}\leq N+1<\eta_{1}$, which yields the second inequality in (6.4). On the other hand, if $k=2^{n}-1$ then by using $\tau_{2^{n}}=1$ it follows that
$\eta_{i+1}\eta_{i+2}=(N+\tau_{2^{n}-1})(N+1+\tau_{2^{n}})\succcurlyeq N(N+2)\succ N(N+1)=\hat{\eta}_{1}\hat{\eta}_{2}.$
So in the following we assume $10\leq k<2^{n}-$ 1. Then by (6.1) we have
$\eta_{i+1}\eta_{i+2}\eta_{i+3}$ $=(N+\tau_{k})(N+1+\tau_{k+1})(N+2-2\tau_{k+1})$
$\succcurlyeq N(N+1)(N+2)\succ N(N+1)(N+1)=\hat{\eta}_{1}\hat{\eta}_{2}\hat{\eta}_{3},$
proving (6.4) when $i=3k+1$.
Case (III). $i=3k+2$ with $0\leq k<2^{n}$. Then $\eta_{i+1}=N+1+\tau_{k+1}\geq N+1>\hat{\eta}_{1}$, which gives the first inequality of (6.4). On the other hand, if $k=2^{n}-1$ then we have $\eta_{i+1}=N+1+\tau_{2^{n}}=N+2=\eta_{1}$; and if $10\leq k<2^{n}-$ 1, then by (6.1) we have
$\eta_{i+1}\eta_{i+2}\eta_{i+3}$ $=(N+1+\tau_{k+1})(N+2-2\tau_{k+1})(N+\tau_{k+1})$
$\preccurlyeq(N+2)N(N+1)\prec(N+2)N(N+2)=\eta_{1}\eta_{2}\eta_{3},$
completing the proof. ∎
Proof of Proposition 6.1.
Note that the Thue-Morse sequence $(\tau_{i})_{i=0}^{\infty}$ satisfies $\tau_{0}=0,\tau_{2k}=\tau_{k}$ and $\tau_{2k+1}=1-\tau_{k}$ for all $k\geq 0$. Then by Definition 1.2 and (6.1) it follows that $(\eta_{i})$ is an expansion of $1$ in base $\beta_{c}=\beta_{c}(M)$. So, by Lemmas 2.1 and 6.2 it follows that $\delta(\beta_{c})=\eta_{1}\eta_{2}\ldots$. Hence, by (6.1) we obtain
$1=\sum_{i=1}^{\infty}\frac{\eta_{i}}{\beta_{c}^{i}}$ $=\sum_{i=0}^{\infty}\frac{N+2-2\tau_{i}}{\beta_{c}^{3i+1}}+\sum_{k=0}^{\infty}\frac{N+\tau_{i}}{\beta_{c}^{3i+2}}+\sum_{i=0}^{\infty}\frac{N+1+\tau_{i+1}}{\beta_{c}^{3i+3}}$
$=\sum_{i=1}^{\infty}\frac{N+1}{\beta_{c}^{i}}+\sum_{i=0}^{\infty}\frac{1}{\beta_{c}^{3i+1}}-\sum_{i=0}^{\infty}\frac{1}{\beta_{c}^{3i+2}}-\frac{2}{\beta_{c}}\sum_{i=0}^{\infty}\frac{\tau_{i}}{\beta_{c}^{3i}}+\frac{1}{\beta_{c}^{2}}\sum_{i=0}^{\infty}\frac{\tau_{i}}{\beta_{c}^{3i}}+\frac{1}{\beta_{c}^{3}}\sum_{i=0}^{\infty}\frac{\tau_{i}}{\beta_{c}^{3i}}$
$=\frac{N+1}{\beta_{c}-1}+\frac{\beta_{c}^{2}-\beta_{c}}{\beta_{c}^{3}-1}+\left(\frac{-2}{\beta_{c}}+\frac{1}{\beta_{c}^{2}}+\frac{1}{\beta_{c}^{3}}\right)\sum_{i=0}^{\infty}\frac{\tau_{i}}{\beta_{c}^{3i}}.$
Since $\beta_{c}>N+2$, we have $\frac{-2}{\beta_{c}}+\frac{1}{\beta_{c}^{2}}+\frac{1}{\beta_{c}^{3}}\neq 0$. Using $\tau_{0}=0$ and rearranging the above equation yields
$\sum_{i=1}^{\infty}\frac{\tau_{i}}{\beta_{c}^{3i}}=\frac{1-\frac{N+1}{\beta_{c}-1}-\frac{\beta_{c}^{2}-\beta_{c}}{\beta_{c}^{3}-1}}{\frac{-2}{\beta_{c}}+\frac{1}{\beta_{c}^{2}}+\frac{1}{\beta_{c}^{3}}}.$
Suppose on the contrary that $\beta_{c}\in(N+2,N+3)$ is algebraic. Then $\beta_{c}^{-3}\in(0,1)$ is also algebraic, and by Lemma 4.4 it follows that the left-hand side of the above equation is transcendental, while the right hand-side is algebraic, leading to a contradiction. So, $\beta_{c}$ is transcendental. ∎
6.1. Patterns in $\widetilde{{\mathbf{U}}}_{\beta_{c},M}$
$x$$y$$\alpha_{00}$$\alpha_{10}$$\alpha_{20}$$\alpha_{30}$$\alpha_{01}$$\alpha_{02}$$\alpha_{03}$$\alpha_{11}$$\alpha_{21}$$\alpha_{12}$$\frac{1}{\beta}$$\frac{2}{\beta}$$\frac{3}{\beta}$$\frac{3}{\beta-1}$$\frac{1}{\beta}$$\frac{2}{\beta}$$\frac{3}{\beta}$$\frac{3}{\beta-1}$
Figure 7. The graph for the first generation of $S_{\beta,M}$ with $M=3$ and $\beta=3$. The convex hull $\Delta_{\beta,M}$ is the triangle with vertices $(0,0),(3/(\beta-1),0)$ and $(0,3/(\beta-1))$. Each light grey triangle corresponds to a $f_{\beta,\alpha_{ij}}(\Delta_{\beta,M})$ for some $\alpha_{ij}\in\Omega_{M}$. Furthermore, the overlap region $O_{\beta,M}$ is the union of eighteen small dark grey triangles.
In the following we will study possible block patterns appearing in sequences of $\widetilde{{\mathbf{U}}}_{\beta_{c},M}$. To describe these patterns we construct a sequence of blocks $\left\{{\mathbf{c}}_{n}\right\}_{n=0}^{\infty}$ in ${\Omega}_{M}^{*}$. Let
$\pazocal A_{3}$ :=$:=\left\{\alpha_{(N+2)(N+1)}\alpha_{N(N+2)}\alpha_{(N+2)N},\quad\alpha_{N(N+1)}\alpha_{(N+1)(N+2)}\alpha_{(N+1)N},\right.$
$\hskip 21.33955pt\left.\alpha_{(N+2)(N+1)}\alpha_{N(N+2)}\alpha_{(N+1)N},\quad\alpha_{N(N+1)}\alpha_{(N+1)(N+2)}\alpha_{(N+2)N}\right\}.$
We define a substitution $\Phi_{3}:\pazocal A_{3}\to\pazocal A_{3}$ by
(6.9) $\begin{split}\alpha_{(N+2)(N+1)}\alpha_{N(N+2)}\alpha_{(N+2)N}\quad&\longleftrightarrow\quad\alpha_{N(N+1)}\alpha_{(N+1)(N+2)}\alpha_{(N+1)N},\\ \alpha_{(N+2)(N+1)}\alpha_{N(N+2)}\alpha_{(N+1)N}\quad&\longleftrightarrow\quad\alpha_{N(N+1)}\alpha_{(N+1)(N+2)}\alpha_{(N+2)N}.\end{split}$
Then the sequence $\left\{{\mathbf{c}}_{n}\right\}_{n=0}^{\infty}\subset\pazocal A_{3}^{*}$ is define recursively by (see Figure 7)
(6.10) ${\mathbf{c}}_{0}=\alpha_{(N+1)(N+1)},\quad{\mathbf{c}}_{1}=\alpha_{(N+2)(N+1)}\alpha_{N(N+2)}\alpha_{(N+1)N},\quad\textrm{and}\quad{\mathbf{c}}_{n+1}={\mathbf{c}}_{n}^{+}\Phi_{3}({\mathbf{c}}_{n}^{+})$
for all $n\geq 1$, where ${\mathbf{c}}_{n}^{+}$ is a word in $\pazocal A_{3}^{*}$ by changing the last digit of ${\mathbf{c}}_{n}$ from $\alpha_{(N+1)N}$ to $\alpha_{(N+2)N}$. Then each ${\mathbf{c}}_{n}$ has length $3\cdot 2^{n}$ and ends with $\alpha_{(N+1)N}$. So,
${\mathbf{c}}_{0}\sim\left(\begin{array}[]{c}N+1\\ N+1\\ N+1\end{array}\right),\quad{\mathbf{c}}_{1}\sim\left(\begin{array}[]{ccc}N+2&N&N+1\\ N+1&N+2&N\\ N&N+1&N+2\end{array}\right),$
and
${\mathbf{c}}_{2}\sim\left(\begin{array}[]{cccccc}N+2&N&N+2&N&N+1&N+1\\ N+1&N+2&N&N+1&N+2&N\\ N&N+1&N+1&N+2&N&N+2\end{array}\right).$
Next we define recursively a sequence of blocks $({\mathbf{v}}_{n})$ which converges to $(\eta_{i})$ componentwisely:
(6.11) ${\mathbf{v}}_{0}=N+1,\quad{\mathbf{v}}_{1}=(N+2)N(N+1),\quad\textrm{and}\quad{\mathbf{v}}_{n+1}:={\mathbf{v}}_{n}^{+}\phi_{3}({\mathbf{v}}_{n}^{+})\quad\forall n\geq 1,$
where $\phi_{3}$ is the block map defined as in (6.2). So, for example,
${\mathbf{v}}_{2}$ $=(N+2)N(N+2)N(N+1)(N+1),$
${\mathbf{v}}_{3}$ $=(N+2)N(N+2)N(N+1)(N+2)\,N(N+1)(N+1)(N+2)N(N+1).$
Then by (6.1) it follows that $(\eta_{i})$ is the componentwise limit of the sequence $({\mathbf{v}}_{n})$, i.e.,
$(\eta_{i})=\lim_{n\to\infty}{\mathbf{v}}_{n}\in\Sigma_{3}^{\mathbb{N}}.$
One can check that $(\eta_{i})$ begins with ${\mathbf{v}}_{n}^{+}$ for any $n\in\mathbb{N}$.
Comparing the definitions of $({\mathbf{v}}_{n})$ and $({\mathbf{c}}_{n})$ in (6.11) and (6.10) respectively, it follows that
(6.12) ${\mathbf{c}}_{n}^{1}={\mathbf{v}}_{n},\quad{\mathbf{c}}_{n}^{2}=((N+1)(N+2)N)^{n}\quad\textrm{and}\quad\overline{{\mathbf{c}}_{n}^{\oplus}}=\phi_{3}({\mathbf{v}}_{n})$
for any $n\geq 0$.
We aim at proving the following pattern result for $\widetilde{{\mathbf{U}}}_{\beta_{c},M}$.
Proposition 6.3.
Let $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{c},M}$ and $n\in\mathbb{N}_{0}$. Then there exists $m_{*}\in\mathbb{N}$ such that for all $m>m_{*}$ the following statements hold.
  1. (i)
    If $d^{1}_{m}<M$ and $d_{m+1}\ldots d_{m+3\cdot 2^{n}}={\mathbf{c}}_{n+1}^{+}$, then $d_{m+3\cdot 2^{n}+1}\ldots d_{m+3\cdot 2^{n+1}}=\Phi_{3}({\mathbf{c}}_{n+1})$ or $\Phi_{3}({\mathbf{c}}_{n+1}^{+}).$
  2. (ii)
    If $d^{1}_{m}>0$ and $d_{m+1}\ldots d_{m+3\cdot 2^{n}}=\Phi_{3}({\mathbf{c}}_{n+1}^{+})$, then $d_{m+3\cdot 2^{n}+1}\ldots d_{m+3\cdot 2^{n+1}}={\mathbf{c}}_{n+1}$ or ${\mathbf{c}}_{n+1}^{+}.$
To prove Proposition 6.3 we need the following lemma.
Lemma 6.4.
For any $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{c},M}\setminus\widetilde{{\mathbf{U}}}_{\beta_{0},M}$ there exists $m_{*}\in\mathbb{N}$ such that
$d_{m_{*}+j}^{1},\quad d_{m_{*}+j^{2}},\quad\overline{d_{m_{*}+j}^{\oplus}}\in\left\{N,N+1,N+2\right\}\quad\forall j\geq 1.$
Proof.
Take $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{c},M}\setminus\widetilde{{\mathbf{U}}}_{\beta_{0},M}$. Note by Proposition 3.1 that $(d_{i}^{1})\in\left\{0,1,\ldots,M\right\}^{\infty}$ does not end with $M^{\infty}$. Then there exists $m_{1}\in\mathbb{N}$ such that $d_{m_{1}}^{1}<M$. Since $\delta(\beta_{c})\preccurlyeq(N+2)^{\infty}$ by Proposition 6.1, we obtain by Proposition 2.3 that $d_{m_{1}+1}^{1}\leq N+2<3N+3=M$; and again we can deduce that $d_{m_{1}+j}^{1}<M$ for all $j\geq 1$. Similarly, since $(d_{i})\notin\widetilde{{\mathbf{U}}}_{\beta_{0},M}$, by Proposition 3.1 there exists $m_{2}>m_{1}$ such that $d_{m_{2}}^{2}<M$. The same argument as above yields that $d_{m_{2}+j}^{2}<M$ for all $j\geq 1$. Also, we can find $m_{*}=m_{3}>m_{2}$ such that $\overline{d_{m_{*}+j}^{\oplus}}<M$ for all $j\geq 1$. So, for all $j\geq 1$ we have
(6.13) $d_{m_{*}+j}^{1}<M,\quad d_{m_{*}+j}^{2}<M\quad\textrm{and}\quad\overline{d_{m_{*}+j}^{\oplus}}<M.$
Since $\delta(\beta_{c})\preccurlyeq(N+2)^{\infty}$, by (2.1) and (6.13) it follows that
$d_{m_{*}+j}^{1},d_{m_{*}+j}^{2},\overline{d_{m_{*}+j}^{\oplus}}\in\left\{N,N+1,N+2\right\}$
for all $j\geq 1$. ∎
First we prove Proposition 6.3 for $n=0$ and $n=1$.
Lemma 6.5.
For any $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{c},M}$ there exists $m_{*}\in\mathbb{N}$ such that for all $m>m_{*}$ the following statements hold.
  1. (i)
    If $d^{1}_{m}<M$ and $d_{m+1}d_{m+2}d_{m+3}={\mathbf{c}}_{1}^{+}$, then $d_{m+4}d_{m+5}d_{m+6}=\Phi_{3}({\mathbf{c}}_{1})$ or $\Phi_{3}({\mathbf{c}}_{1}^{+}).$
  2. (ii)
    If $d^{1}_{m}>0$ and $d_{m+1}d_{m+2}d_{m+3}=\Phi_{3}({\mathbf{c}}_{1}^{+})$, then $d_{m+4}d_{m+5}d_{m+6}={\mathbf{c}}_{1}$ or ${\mathbf{c}}_{1}^{+}.$
Proof.
Since the proof of (ii) is similar, we only prove (i). Take $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{c},M}$. If $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{0},M}$, then by Proposition 3.1 it follows that $d_{i+1}d_{i+2}d_{i+3}\notin\left\{{\mathbf{c}}_{1}^{+},\Phi_{3}({\mathbf{c}}_{1}^{+})\right\}$ for all $i\geq 0$, and we are done. In the following we assume $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{c},M}\setminus\widetilde{{\mathbf{U}}}_{\beta_{0},M}$. Then by Lemma 6.4 there exists $m_{*}\in\mathbb{N}$ such that $d_{m_{*}+j}^{1},d_{m_{*}+j}^{2},\overline{d_{m_{*}+j}^{\oplus}}\in\left\{N,N+1,N+2\right\}$ for all $j\geq 1$. Take $m>m_{*}$. Suppose $d_{m}^{1}<M$ and $d_{m+1}d_{m+2}d_{m+3}={\mathbf{c}}_{1}^{+}=\alpha_{(N+2)(N+1)}\alpha_{N(N+2)}\alpha_{(N+2)N}$. Then by Propositions 2.3 and 6.1 it follows that $d_{m+4}\in\left\{\alpha_{N(N+1)},\alpha_{N(N+2)}\right\}$.
If $d_{m+4}=\alpha_{N(N+2)}$, then by Propositions 2.3 and 6.1 we obtain that
$d_{m+1}\ldots d_{m+9}\sim\begin{pmatrix}(N+2)N(N+2)&N(N+1)(N+2)&N(N+1)(N+1)\\ (N+1)(N+2)N&(N+2)N(N+1)&(N+2)N(N+1)\\ N(N+1)(N+1)&(N+1)(N+2)N&(N+1)(N+2)(N+1)\end{pmatrix}$
or
$d_{m+1}\ldots d_{m+9}\sim\begin{pmatrix}(N+2)N(N+2)&N(N+1)(N+2)&N(N+1)(N+1)\\ (N+1)(N+2)N&(N+2)N(N+1)&(N+1)(N+2)(N+1)\\ N(N+1)(N+1)&(N+1)(N+2)N&(N+2)N(N+1)\end{pmatrix}.$
However, the forbidden block $(N+1)(N+2)(N+1)$ occurs in $\overline{d_{m+1}^{\oplus}}\ldots\overline{d_{m+9}^{\oplus}}$ in the first cases, and occurs in $d_{m+1}^{2}\ldots d_{m+9}^{2}$ in the second case, leading to a contradiction with $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{c},M}$. So, $d_{m+4}=\alpha_{N(N+1)}$.
By (2.1), Propositions 2.3 and 6.1 it follows that $d_{m+5}=\alpha_{(N+1)(N+2)}$ and $d_{m+6}=\alpha_{(N+2)N}$ or $\alpha_{(N+1)N}$. This implies that $d_{m+4}d_{m+5}d_{m+6}=\Phi_{3}({\mathbf{c}}_{1})$ or $\Phi_{3}({\mathbf{c}}_{1}^{+})$ as desired. ∎
Lemma 6.6.
For any $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{c},M}$ there exists $m_{*}\in\mathbb{N}$ such that for all $m>m_{*}$ the following statements hold.
  1. (i)
    If $d^{1}_{m}<M$ and $d_{m+1}\ldots d_{m+6}={\mathbf{c}}_{2}^{+}$, then $d_{m+7}\ldots d_{m+12}=\Phi_{3}({\mathbf{c}}_{2})$ or $\Phi_{3}({\mathbf{c}}_{2}^{+}).$
  2. (ii)
    If $d^{1}_{m}>0$ and $d_{m+1}\ldots d_{m+6}=\Phi_{3}({\mathbf{c}}_{2}^{+})$, then $d_{m+7}\ldots d_{m+12}={\mathbf{c}}_{2}$ or ${\mathbf{c}}_{2}^{+}.$
Proof.
Since the proof of (ii) is similar, here we only prove (i). Take $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{c},M}$. By the same argument as in the proof of Lemma 6.5 we may assume $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{c},M}\setminus\widetilde{{\mathbf{U}}}_{\beta_{0},M}$. Then by Lemma 6.4 there exists $m_{*}\in\mathbb{N}$ such that each component of $d_{m_{*}+j}$ belongs to $\left\{N,N+1,N+2\right\}$ for all $j\geq 1$. Take $m>m_{*}$. Suppose $d_{m}^{1}<M$ and $d_{m+1}\ldots d_{m+6}={\mathbf{c}}_{2}^{+}=\alpha_{(N+2)(N+1)}\alpha_{N(N+2)}\alpha_{(N+2)(N+1)}\alpha_{N(N+1)}\alpha_{(N+1)(N+2)}\alpha_{(N+2)N}$ for some $m>m_{*}$. Then by (2.1), Propositions 2.3 and 6.1 it follows that $d_{m+7}\in\left\{\alpha_{N(N+1)},\alpha_{N(N+2)}\right\}$.
If $d_{m+7}=\alpha_{N(N+2)}$, then by (2.1), Propositions 2.3 and 6.1 we obtain
$d_{m+1}\ldots d_{m+9}\sim\begin{pmatrix}(N+2)N(N+2)&N(N+1)(N+2)&N(N+1)(N+1)\\ (N+1)(N+2)N&(N+1)(N+2)N&(N+2)N(N+1)\\ N(N+1)(N+1)&(N+2)N(N+1)&(N+1)(N+2)(N+1)\end{pmatrix}.$
Thus, the forbidden block $(N+1)(N+2)(N+1)$ appears in $\overline{d_{m+1}^{\oplus}}\ldots\overline{d_{m+9}^{\oplus}}$, which leads to a contradiction with $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{c},M}$. So, $d_{m+7}=\alpha_{N(N+1)}$.
Hence, by (2.1), Propositions 2.3 and 6.1 it follows that
$d_{m+7}d_{m+8}d_{m+9}=\alpha_{N(N+1)}\alpha_{(N+1)(N+2)}\alpha_{(N+1)N}=\Phi_{3}({\mathbf{c}}_{1}^{+}).$
Since $d_{m+6}^{1}=N+2>0$, by Lemma 6.5 (ii) we obtain $d_{m+10}d_{m+11}d_{m+12}={\mathbf{c}}_{1}$ or ${\mathbf{c}}_{1}^{+}$. So,
$d_{m+7}\ldots d_{m+12}=\Phi_{3}({\mathbf{c}}_{1}^{+}){\mathbf{c}}_{1}=\Phi_{3}({\mathbf{c}}_{2}^{+})\quad\textrm{or}\quad\Phi_{3}({\mathbf{c}}_{1}^{+}){\mathbf{c}}_{1}^{+}=\Phi_{3}({\mathbf{c}}_{2}),$
completing the proof. ∎
Proof of Proposition 6.3.
Since the proof of (ii) is similar, we only prove (i). Take $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{c},M}$. By Proposition 3.1 it suffices to consider $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{c},M}\setminus\widetilde{{\mathbf{U}}}_{\beta_{0},M}$. In view of Lemma 6.4, there exists $m_{*}\in\mathbb{N}$ such that each component of $d_{m_{*}+j}$ belongs to $\left\{N,N+1,N+2\right\}$ for all $j\geq 1$. Take $m>m_{*}$. Suppose $d_{m}^{1}<M$ and $d_{m+1}\ldots d_{m+3\cdot 2^{n}}={\mathbf{c}}_{n+1}^{+}$ for some $m>m_{*}$. We will prove
(6.14) $d_{m+3\cdot 2^{n}+1}\ldots d_{m+3\cdot 2^{n+1}}=\Phi_{3}({\mathbf{c}}_{n+1}^{+})\quad\textrm{or}\quad\Phi_{3}({\mathbf{c}}_{n+1}),$
which will be done by induction on $n$. If $n=0$ or $1$, then (6.14) holds by Lemmas 6.5 and 6.6. Now suppose (6.14) holds for all $n\leq k$ for some $k\in\mathbb{N}$, and we consider it for $n=k+1$.
Note that $d_{m+1}\ldots d_{m+3\cdot 2^{k+1}}={\mathbf{c}}_{k+2}^{+}$. By (6.11) and (6.12) it follows that
$d_{m+1}^{1}\ldots d_{m+3\cdot 2^{k+1}}^{1}={\mathbf{v}}_{k+2}^{+}={\mathbf{v}}_{k+1}^{+}\phi_{3}({\mathbf{v}}_{k+1}).$
Since by (6.1) the sequence $\delta(\beta_{c})$ begins with ${\mathbf{v}}_{k+2}^{+}\phi_{3}({\mathbf{v}}_{k+2})$, by (6.11) and Proposition 2.3 it follows that
(6.15) $d_{m+3\cdot 2^{k+1}+1}^{1}\ldots d_{m+3\cdot 2^{k+2}}^{1}\preccurlyeq\phi_{3}({\mathbf{v}}_{k+2})=\phi_{3}({\mathbf{v}}_{k+1}^{+}){\mathbf{v}}_{k+1}^{+}=\phi_{3}({\mathbf{v}}_{k}^{+}){\mathbf{v}}_{k}{\mathbf{v}}_{k+1}^{+}.$
Note that $d_{m+1}\ldots d_{m+3\cdot 2^{k+1}}={\mathbf{c}}_{k+2}^{+}={\mathbf{c}}_{k+1}^{+}\Phi_{3}({\mathbf{c}}_{k}^{+}){\mathbf{c}}_{k}^{+}$. Then $d_{m+3\cdot(2^{k}+2^{k-1})}^{1}=N+1<M$ and $d_{m+3(2^{k}+2^{k-1})+1}\ldots d_{m+3\cdot 2^{k+1}}={\mathbf{c}}_{k}^{+}$. Then by the induction hypothesis for $n=k-1$ we have
$d_{m+3\cdot 2^{k+1}+1}\ldots d_{m+3\cdot(2^{k+1}+2^{k-1})}=\Phi_{3}({\mathbf{c}}_{k}^{+})\quad\textrm{or}\quad\Phi_{3}({\mathbf{c}}_{k}).$
If it takes $\Phi_{3}({\mathbf{c}}_{k})$, then $d^{1}_{m+3\cdot 2^{k+1}+1}\ldots d^{1}_{m+3\cdot(2^{k+1}+2^{k-1})}=\phi_{3}({\mathbf{v}}_{k})\succ\phi_{3}({\mathbf{v}}_{k}^{+})$, leading to a contradiction with (6.15). So, $d_{m+3\cdot 2^{k+1}+1}\ldots d_{m+3\cdot(2^{k+1}+2^{k-1})}=\Phi_{3}({\mathbf{c}}_{k}^{+})$. Since $d_{m+3\cdot 2^{k+1}}^{1}=N+2>0$, by the induction hypothesis we obtain that $d_{m+3\cdot 2^{k+1}+1}\ldots d_{m+3\cdot(2^{k+1}+2^{k})}=\Phi_{3}({\mathbf{c}}_{k+1}^{+})$ or $\Phi_{3}({\mathbf{c}}_{k+1})$. In terms of (6.15) the block $\Phi_{3}({\mathbf{c}}_{k+1})$ is not allowed. So,
$d_{m+3\cdot 2^{k+1}+1}\ldots d_{m+3\cdot(2^{k+1}+2^{k})}=\Phi_{3}({\mathbf{c}}_{k+1}^{+}).$
Again by using $d_{m+3\cdot 2^{k+1}}^{1}=N+2>0$ and the induction hypothesis it follows that
$d_{m+3\cdot 2^{k+1}+1}\ldots d_{m+3\cdot 2^{k+2}}=\Phi_{3}({\mathbf{c}}_{k+2}^{+})\quad\textrm{or}\quad\Phi_{3}({\mathbf{c}}_{k+2}),$
proving (6.14) for $n=k+1$. This completes the proof by induction. ∎
6.2. Phase transition of $\widetilde{{\mathbf{U}}}_{\beta,M}$ for $M=3N+3$
Note by (6.11) that ${\mathbf{v}}_{n}=\eta_{1}\ldots\eta_{3\cdot 2^{n-1}}^{-}$ for all $n\geq 1$. Then by Lemma 6.2 it follows that $\sigma^{j}({\mathbf{v}}_{n}^{\infty})\preccurlyeq{\mathbf{v}}_{n}^{\infty}$ for all $j\geq 0$. So, by Lemma 2.1 there exists $\beta_{n}=\beta_{n}(M)\in(N+2,N+3)$ such that
$\delta(\beta_{n})={\mathbf{v}}_{n}^{\infty}.$
Since ${\mathbf{v}}_{n}\nearrow(\eta_{i})=\delta(\beta_{c})$, again by Lemma 2.1 it follows that
$\beta_{G}=\beta_{0}<\beta_{1}<\cdots<\beta_{n}<\beta_{n+1}<\cdots,\quad\textrm{and}\quad\beta_{n}\nearrow\beta_{c}\quad\textrm{as }n\to\infty.$
Proposition 6.7.
Let $M=3N+3$ and $n\in\mathbb{N}_{0}$. Then any sequence in $\widetilde{{\mathbf{U}}}_{\beta_{n+1},M}\setminus\widetilde{{\mathbf{U}}}_{\beta_{n},M}$ must end in
$\bigcup_{j=0}^{2}\left\{\theta^{j}({\mathbf{c}}_{n}^{\infty}),\theta^{j}(\Phi_{3}({\mathbf{c}}_{n})^{\infty})\right\},$
where $\theta$ is the cyclic permutation defined in (2.2).
We will prove Proposition 6.7 by induction on $n$. First we prove it for $n=0,1$.
Lemma 6.8.
The following statements hold.
  1. (i)
    Any sequence in $\widetilde{{\mathbf{U}}}_{\beta_{1},M}\setminus\widetilde{{\mathbf{U}}}_{\beta_{0},M}$ must end in ${\mathbf{c}}_{0}^{\infty}$;
  2. (ii)
    Any sequence in $\widetilde{{\mathbf{U}}}_{\beta_{2},M}\setminus\widetilde{{\mathbf{U}}}_{\beta_{1},M}$ must end in $\bigcup_{j=0}^{2}\left\{\theta^{j}({\mathbf{c}}_{1}^{\infty}),\theta^{j}(\Phi_{3}({\mathbf{c}}_{1})^{\infty})\right\}$.
Proof.
For (i) we take $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{1},M}\setminus\widetilde{{\mathbf{U}}}_{\beta_{0},M}$. Note that $\delta(\beta_{1})={\mathbf{v}}_{1}^{\infty}=((N+2)N(N+1))^{\infty}$. By the same argument as in the proof of Lemma 6.4 there exists $m\in\mathbb{N}$ such that each component of $d_{m+j}$ belongs to $\left\{N,N+1,N+2\right\}$ for all $j\geq 1$. Then by (2.1) it follows that
$d_{m+1}=\alpha_{(N+1)(N+1)}\quad\textrm{or}\quad d_{m+1}\in\bigcup_{j=0}^{2}\left\{\theta^{j}(\alpha_{(N+2)(N+1)}),\theta^{j}(\alpha_{(N+2)N})\right\}.$
Suppose $d_{m+1}=\alpha_{(N+2)(N+1)}$. By Proposition 2.3 and using $\delta(\beta_{1})=((N+2)N(N+1))^{\infty}$ it follows that
$d_{m+1}d_{m+2}\ldots=(\alpha_{(N+2)(N+1)}\alpha_{N(N+2)}\alpha_{(N+1)N})^{\infty}\quad\textrm{or}\quad(\alpha_{(N+2)(N+1)}\alpha_{N(N+1)}\alpha_{(N+1)(N+2)})^{\infty}.$
However, in either case the component sequence $d_{m+1}^{1}d_{m+2}^{1}\ldots$ ends with $((N+2)N(N+1))^{\infty}={\mathbf{v}}_{1}^{\infty}$, which leads to a contradiction with $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{1},M}$. Similarly, if $d_{m+1}$ equals any other digit in $\bigcup_{j=0}^{2}\left\{\theta^{j}(\alpha_{(N+2)(N+1)}),\theta^{j}(\alpha_{(N+2)N})\right\}$, we can deduce the same contradiction. So, $d_{m+1}=\alpha_{(N+1)(N+1)}$; and therefore by the same argument as above we can prove that $d_{m+1}d_{m+2}\ldots=\alpha_{(N+1)(N+1)}^{\infty}={\mathbf{c}}_{0}^{\infty}$.
Next we prove (ii). Take $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{2},M}\setminus\widetilde{{\mathbf{U}}}_{\beta_{1},M}$. Note that
(6.16) $\delta(\beta_{2})={\mathbf{v}}_{2}^{\infty}=((N+2)N(N+2)N(N+1)(N+1))^{\infty}.$
By the same argument as in the proof of Lemma 6.4 there exists $m_{*}\in\mathbb{N}$ such that each component of $d_{m_{*}+j}$ belongs to $\left\{N,N+1,N+2\right\}$ for all $j\geq 1$. Again by (i) there exists $m>m_{*}$ such that
$d_{m+j}\in\bigcup_{k=0}^{2}\left\{\theta^{k}(\alpha_{(N+2)(N+1)}),\theta^{k}(\alpha_{(N+2)N})\right\}.$
Without loss of generality we may assume $d_{m+1}=\alpha_{(N+2)(N+1)}$. Then by (6.16) and Proposition 2.3 it follows that
(6.17) $d_{m+1}d_{m+2}d_{m+3}\in\left\{\alpha_{(N+2)(N+1)}\alpha_{N(N+2)}\alpha_{(N+1)N},\alpha_{(N+2)(N+1)}\alpha_{N(N+1)}\alpha_{(N+1)(N+2)}\right\}$
or
(6.18) $d_{m+1}d_{m+2}d_{m+3}\in\left\{\alpha_{(N+2)(N+1)}\alpha_{N(N+2)}\alpha_{(N+2)N},\alpha_{(N+2)(N+1)}\alpha_{N(N+1)}\alpha_{(N+2)(N+1)}\right\}.$
We will prove by contradiction that (6.18) never occurs.
If $d_{m+1}d_{m+2}d_{m+3}=\alpha_{(N+2)(N+1)}\alpha_{N(N+2)}\alpha_{(N+2)N}$, then by (6.16) and Proposition 2.3 it follows that
$d_{m+1}d_{m+2}\ldots=(\alpha_{(N+2)(N+1)}\alpha_{N(N+2)}\alpha_{(N+2)N}\alpha_{N(N+1)}\alpha_{(N+1)(N+2)}\alpha_{(N+1)N})^{\infty},$
which implies that $d_{m+1}^{1}d_{m+2}^{1}\ldots={\mathbf{v}}_{2}^{\infty}=\delta(\beta_{2})$, leading to a contradiction with $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{2},M}$. If $d_{m+1}d_{m+2}d_{m+3}=\alpha_{(N+2)(N+1)}\alpha_{N(N+1)}\alpha_{(N+2)(N+1)}$, then by (6.16) and Proposition 2.3 we obtain that
$d_{m+4}d_{m+5}\ldots=(\alpha_{N(N+2)}\alpha_{(N+1)N}\alpha_{(N+1)(N+2)}\alpha_{(N+2)N}\alpha_{N(N+1)}\alpha_{(N+2)(N+1)})^{\infty},$
which implies that $d_{m+4}^{1}d_{m+5}^{1}\ldots$ ends with ${\mathbf{v}}_{2}^{\infty}=\delta(\beta_{2})$, again leading to a contradiction with $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{2},M}$. So, we must have (6.17).
By the same argument as above we can even prove that any block $d_{m+j}d_{m+j+1}d_{m+j+2}$ can not belong to
$\bigcup_{k=0}^{2}\left\{\theta^{k}(\alpha_{(N+2)(N+1)}\alpha_{N(N+2)}\alpha_{(N+2)N}),\theta^{k}(\alpha_{(N+2)(N+1)}\alpha_{N(N+1)}\alpha_{(N+2)(N+1)})\right\}.$
From this we can deduce that $d_{m+1}d_{m+2}\ldots$ must end with ${\mathbf{c}}_{1}^{\infty}$ as required.
If $d_{m+1}$ is taken some other digit from $\bigcup_{k=0}^{2}\left\{\theta^{k}(\alpha_{(N+2)(N+1)}),\theta^{k}(\alpha_{(N+2)N})\right\}$, we can prove by the same argument as above that $d_{m+1}d_{m+2}\ldots$ ends with $\theta^{k}({\mathbf{c}}_{1}^{\infty})$ for some $k=0,1,2$. This completes the proof. ∎
Proof of Proposition 6.7.
We will prove the result by induction on $n$. By Lemma 6.8 the result holds for $n=0,1$. Now suppose it holds for all $n\leq k$ for some $k\geq 1$, and we consider it for $n=k+1$. Take $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{k+2},M}\setminus\widetilde{{\mathbf{U}}}_{\beta_{k+1},M}$. By Lemma 6.4 there exists $m_{*}\in\mathbb{N}$ such that each component of $d_{m_{*}+j}$ belongs to $\left\{N,N+1,N+2\right\}$ for all $j\geq 1$. Note that $\widetilde{{\mathbf{U}}}_{\beta_{k+1},M}=\widetilde{{\mathbf{U}}}_{\beta_{0},M}\cup\bigcup_{j=0}^{k}(\widetilde{{\mathbf{U}}}_{\beta_{j+1},M}\setminus\widetilde{{\mathbf{U}}}_{\beta_{j},M})$. By Proposition 3.1 and the induction hypothesis it follows that any sequence in $\widetilde{{\mathbf{U}}}_{\beta_{k+1},M}$ must end in
(6.19) $\bigcup_{j=0}^{k}\bigcup_{l=0}^{2}\left\{\theta^{l}({\mathbf{c}}_{j}^{\infty}),\theta^{l}(\Phi_{3}({\mathbf{c}}_{j}^{\infty}))\right\}.$
Observe by (6.12) that ${\mathbf{c}}_{j}^{1}={\mathbf{v}}_{j}$ for all $j\geq 0$, and ${\mathbf{v}}_{0}^{\infty}\prec{\mathbf{v}}_{1}^{\infty}\prec\cdots\prec{\mathbf{v}}_{k}^{\infty}\prec{\mathbf{v}}_{k+1}^{\infty}$. Since $\delta(\beta_{j})={\mathbf{v}}_{j}^{\infty}$ for all $j\geq 0$, by (6.10) and (6.19) there exists $m>m_{*}$ such that $d_{m+1}\ldots d_{m+3\cdot 2^{k-1}}\in\bigcup_{j=0}^{2}\left\{\theta^{j}({\mathbf{c}}_{k}^{+}),\theta^{j}(\Phi_{3}({\mathbf{c}}_{k}^{+}))\right\}$. Without loss of generality we may assume $d_{m+1}\ldots d_{m+3\cdot 2^{k}}={\mathbf{c}}_{k}^{+}$. Note that $d^{1}_{m}\leq N+2<M$. Then by Proposition 6.3 it follows that
(6.20) $d_{m+1}\ldots d_{m+3\cdot 2^{k}}={\mathbf{c}}_{k+1}\quad\textrm{or}\quad{\mathbf{c}}_{k+1}^{+}.$
If $d_{m+1}\ldots d_{m+3\cdot 2^{k}}={\mathbf{c}}_{k+1}^{+}$, the by Proposition 6.3 and (6.12) it follows that
$d_{m+1}^{1}d_{m+2}^{1}\ldots\succcurlyeq{\mathbf{v}}_{k+2}^{\infty}=\delta(\beta_{k+2}),$
which leads to a contradiction with $(d_{i})\in\widetilde{{\mathbf{U}}}_{\beta_{k+2},M}$.
Hence, by (6.20) we must have $d_{m+1}\ldots d_{m+3\cdot 2^{k}}={\mathbf{c}}_{k+1}={\mathbf{c}}_{k}^{+}\Phi_{3}({\mathbf{c}}_{k}^{+})$. Again, by Proposition 6.3 we obtain that $d_{m+3\cdot 2^{k}+1}\ldots d_{m+3(2^{k}+2^{k-1})}={\mathbf{c}}_{k}$ or ${\mathbf{c}}_{k}^{+}.$ If it takes ${\mathbf{c}}_{k}$, then $d_{m+3\cdot 2^{k-1}+1}\ldots d_{m+3\cdot(2^{k}+2^{k-1})}=\Phi_{3}({\mathbf{c}}_{k}^{+}){\mathbf{c}}_{k}=\Phi_{3}({\mathbf{c}}_{k+1}^{+})$, and the same argument as above will lead to a contradiction. So, $d_{m+3\cdot 2^{k}+1}\ldots d_{m+3(2^{k}+2^{k-1})}={\mathbf{c}}_{k}^{+}$. Continuing this argument we can prove that $d_{m+1}d_{m+2}\ldots=({\mathbf{c}}_{k}^{+}\Phi_{3}({\mathbf{c}}_{k}^{+}))^{\infty}={\mathbf{c}}_{k+1}^{\infty},$ which establishes the result for $n=k+1$. This completes the proof by induction. ∎
Proof of Theorem 1.4 for $M=3N+3$.
The proof is similar to that for $M=3N+1$. By using Proposition 6.7 we can deduce that for $\beta<\beta_{c}(M)$ the set $\widetilde{{\mathbf{U}}}_{\beta,M}$ is at most countable; and by Proposition 6.3 we can prove that for $\beta=\beta_{c}(M)$ the set $\widetilde{{\mathbf{U}}}_{\beta,M}$ is uncountable but has zero Hausdorff dimension. Furthermore, for $\beta>\beta_{c}(M)$ we can construct a subshift of finite type $Z_{k}\subset\widetilde{{\mathbf{U}}}_{\beta,M}$ having positive Hausdorff dimension, where $Z_{k}$ is analogous to $X_{k}$ as described in Figure 3 with $\mathbf{a}_{k},\mathbf{a}_{k}^{+},\Phi_{1}(\mathbf{a}_{k})$ and $\Phi_{1}(\mathbf{a}_{k}^{+})$ replaced by ${\mathbf{c}}_{k},{\mathbf{c}}_{k}^{+},\Phi_{3}({\mathbf{c}}_{k})$ and $\Phi_{3}({\mathbf{c}}_{k}^{+})$, respectively. ∎
Acknowledgements
The authors thank Wolfgang Steiner for his insight on an alternate definition of the Thue-Morse type sequences. The first author was supported by NSFC No. 12301108. The second author was supported by Chongqing NSF: CQYC20220511052. The third author was supported by NSFC No. 12071148 and 12471085.